### v1

# Leveraging Discrete Function Decomposability for Scientific Design

Abstract
In the era of AI-driven science and engineering, we often want to design discrete objects in silico according to user-specified properties. For example, we may wish to design a protein to bind its target, arrange components within a circuit to minimize latency, or find materials with certain properties. Given a property predictive model, in silico design typically involves training a generative model over the design space (e.g., protein sequence space) to concentrate on designs with the desired properties. Distributional optimization—which can be formalized as an estimation of distribution algorithm or as reinforcement learning policy optimization—finds the generative model that maximizes an objective function in expectation. Optimizing a distribution over discrete-valued designs is in general challenging because of the combinatorial nature of the design space. However, many property predictors in scientific applications are decomposable in the sense that they can be factorized over design variables in a way that could in principle enable more effective optimization. For example, amino acids at a catalytic site of a protein may only loosely interact with amino acids of the rest of the protein to achieve maximal catalytic activity. Current distributional optimization algorithms are unable to make use of such decomposability structure. Herein, we propose and demonstrate use of a new distributional optimization algorithm, Decomposition-Aware Distributional Optimization (DADO), that can leverage any decomposability defined by a junction tree on the design variables, to make optimization more efficient. At its core, DADO employs a soft-factorized “search distribution”—a learned generative model—for efficient navigation of the search space, invoking graph message-passing to coordinate optimization across linked factors.
1 Design in discrete state spaces
Figure 1: Key components of DADO. a, DADO requires as input an objective function in its decomposed form, $f(x)=C_{1}(\hat{x}_{1})+C_{2}(\hat{x}_{2}),\dots,C_{\kappa}(\hat{x}_{\kappa})$, which corresponds to a junction tree. Here we show a junction tree with nodes of size 1, i.e., a regular tree, for simplicity. Variables with edges interact together to directly influence $f$. Some variables participate in multiple component functions, requiring coordination in the form of message-passing. b, To update the search distribution at each iteration, naive EDAs weight entire samples drawn from a joint distribution over all design variables by scoring with $f(x)$. Shade of green denotes more to less optimal scores. In contrast, DADO leverages the decomposition of $f(x)$ to weight samples in a more local manner, according to the decomposition. Specifically, DADO uses message-passing to compute value functions that account for $x_{i}$ interacting with its descendants. Correspondingly, the value functions serve as the weights for each part of the search distribution, which is factorized like $f$. Optional shaping function $W$ is omitted for clarity. c, Example performance comparison on a synthetic problem with an exact tree decomposition over a discrete design space of size $20^{50}$ ($D=20,L=50$). Each of the two methods drew 100 samples per iteration. We evaluated these samples with $f(x)$, computing the per-iteration mean and 95% confidence interval. Results shown were averaged over 20 random seeds for the same $f(x)$ (details in Sec. 4.1, Fig. 2). The p-value shown is from a two-sided paired t-test that AUC of the per-iteration means is different between methods, using the 20 mean curves.
The integration of AI into scientific research has opened new avenues for property-driven, in silico design of discrete objects—from molecular structures like proteins, to engineered systems like circuits—where computational methods guide design with user-specified properties. For example, we may aim to design an amino acid sequence for a protein so that the protein binds its target.
Given a property predictive model, $f(x)$, the simplest version of in silico design entails enumerating all possible designs, $x=[x_{1},x_{2},...,x_{L}]\in\mathcal{X}$ (e.g., all possible amino acids of length $L$), evaluating each one under the predictive model for property, $s=f(x)$, and choosing the design with the highest $s$. Such a setup is complicated by two primary challenges. First, in most realistic problems, the design space is too large to fully enumerate. Second, if $f(x)$ has parameters estimated from data, then it is most likely not accurate over the whole space. While much work has attempted to address the second problem (e.g.brookes19_cbas; trabucco21_coms; uehara24_braid), little attention has been placed on the first one in the context of decomposable predictive models. Herein, we focus on this issue: performing efficient optimization in high-dimensional discrete design spaces that can be decomposed according to structure in $f(x)$. As the second problem is complementary to ours, we do not here consider it—solutions for each can be implemented together. We focus our examples and experiments on designing amino acid sequences, but our proposed method is general, applicable to design on any discrete space. For a $D$-amino acid alphabet, where typically $D=20$, finding the highest-scoring design, $x^{*}$, naively requires $D^{L}$ evaluations of $f(x)$. To develop some intuition, in this paragraph only, assume that $D=2$ so that $x$ is binary. First, consider the simple but unrealistic case of full linear additivity, where $f(x)=\beta_{1}x_{1}+\beta_{2}x_{2},\ldots,\beta_{L}x_{L}$, for scalar parameters $\{\beta_{i}\}$. In such a scenario, one can find $x^{*}$ by optimizing each of the $L$ components of $x$ separately because they do not interact with each other at all, requiring evaluating only $D\times L\ll D^{L}$ partial designs. A more realistic setting would allow for more complicated functional forms while still having some notion of linear additivity. For example, consider the form $f(x)=C_{1}(\hat{x}_{1})+C_{2}(\hat{x}_{2}),\dots,C_{\kappa}(\hat{x}_{\kappa})$, where $C_{i}$ denotes an arbitrary “component” function on a set of design variables, $\hat{x}_{i}$ (e.g., $\hat{x}_{i}=\{x_{1},x_{3},x_{8}\}$), such as $C_{1}(\hat{x}_{1})=C_{1}(x_{1},x_{3},x_{8})=\exp(7x_{1}x_{3}-3x_{3}x_{8})$. As long as sets of variables don’t overlap, i.e., $x_{i}$ only appears in a single component function $C_{j}$, we can separately optimize each component $C_{j}$, which is over fewer variables than $f(x)$, resulting in an easier and more efficient optimization. In the most general setting, encompassing any function on discrete variables, sets of variables can overlap across component functions, thereby tying them together and making the optimization correspondingly harder—but as we will see, still much easier than naive optimization of $f$ in many cases (Fig. 1). This difference in optimization difficulty between the decomposed and naive scenarios increases with the alphabet size and number of design variables, which are both large in many realistic settings. If $f$ has no decomposability at all, it can be represented as a single component function over all variables, namely, itself, in which case there is no efficiency to be gained. However, in most real problems, we expect some level of decomposability because the world is generally structured. For example, in designing a protein to bind to a target, a few key sequence positions may make up the binding interface, which jointly, primarily dictate the binding strength, whereas the other positions may be involved, largely separately from the binding site, in stabilizing the protein for binding. Our approach will not require full decomposability so that these types of factors can still interact with each other. We refer to this as “soft-decomposability”, whereby some variables can participate in multiple component functions.
To optimize a soft-decomposable function, one can employ classical message-passing algorithms, which leverage decomposability to exactly find the global optimum with the lowest possible time complexity (vlassis04_maxplus). However, for reasons discussed momentarily, we are interested in distributional optimization (DO), wherein a standard optimization problem over the design space, $x^{*}=\arg\max_{x\in\mathcal{X}}f(x)$, is replaced by one over the parameters of a generative model, $p_{\theta}(x)$, namely, $\theta^{*}=\arg\max_{\theta}\mathbb{E}_{p_{\theta}(x)}\left[f(x)\right]$. These two optimization problems are equivalent in that the optimum of each is the same when $p_{\theta}(x)$ has capacity to place all of its mass on the maximal designs. Instead of naive, decomposition-unaware DO, we will introduce a decomposition-aware version of DO that takes advantage of soft-decomposability in $f$ by way of a soft-factorized $p_{\theta}(x)$, each factor of which is responsible for a single component of $x$ (e.g., $\hat{x}_{i}$). This factorization results in an overall effective search space that is substantially smaller than $\mathcal{X}$, like in our examples above. Moreover, we will show that each factor distribution can be updated separately from all others, such that given a fixed sampling budget for training $p_{\theta}(x)$, each factor can apply the full budget to its smaller subset of design dimensions. As a result, our decomposition-aware DO algorithm will be more statistically efficient than naive DO and thus navigate toward promising regions of $\mathcal{X}$ more quickly.
Our interest in the distributional optimization formulation is motivated by its extensibility. First, DO enables one to directly use innovations from the Estimation of Distribution Algorithm (EDA) (larranaga01_eda; brookes20_eda_em) and policy optimization (peters07_rwr; peng19_awr) communities. Of particular note are methods that enable combining a pre-trained, unconditional generative model, $p(x)$, with a property predictor, $p(y|x)$, to execute Bayes rule so as to obtain a sampling distribution that can be used for design, $p(x|y\in Y)$ (e.g.brookes19_cbas; Fannjiang2020; uehara24_braid). Second, as the key object that navigates the search space, $p_{\theta}(x)$, is a generative model, we stand to benefit from advances in generative modeling. Herein, for clarity of contribution, we focus on the purest form of the EDA, without entropy regularization, a prior, or conditioning, leaving such extensions to future work.
Contributions.
We develop a DO algorithm in the form of a generalized EDA/policy optimization algorithm that can leverage any decomposability in $f(x)$ in the form described earlier as $f(x)=C_{1}(\hat{x}_{1})+C_{2}(\hat{x}_{2}),\dots,C_{\kappa}(\hat{x}_{\kappa})$ (also equivalently described by a junction tree), to more efficiently navigate the design space and find designs with high $f(x)$ (Fig. 1). We call our method Decomposition-Aware Distributional Optimization (DADO). We first empirically investigate DADO on entirely synthetic problems, illustrating that DADO indeed yields designs with high $f(x)$ more quickly than a version that is not decomposition-aware. Next we investigate DADO for problems anchored on real protein data, further substantiating that leveraging decomposition helps.
As DADO requires as input a soft-decomposed objective, $f(x)$, for our protein problems we construct these as follows. We first obtain a graph topology of which variables (protein positions) are connected by thresholding residue distances from an AlphaFold3 structure, and adding an edge between pairs of positions within this threshold (abramson24_af3; brookes22_sparsity). From this graph we can construct a junction tree, which in turn directly maps to the functional form of the predictive model. Recall that junction trees represent arbitrarily complex relationships between variables by transforming any graph on those variables into a tree structure, where each node contains a set of the original variables, and where these sets may be overlapping in adjacent nodes (lauritzen88_JT). Given the junction tree topology, we can then fit a neural network adherent to the topology on sequence-property pairs, $(x^{k},y^{k})$. More precisely, each component $C_{i}(\hat{x}_{i})$ is parameterized and the sum of all components, $f(x^{k})=\sum_{i}^{\kappa}C_{i}(\hat{x}^{k}_{i})$, regressed onto property labels $y^{k}$. Analogous procedures in different application areas could include the following, for example. In circuit design, certain topologies of components may be prescribed ahead of time according to production constraints or domain knowledge. When designing a telescope, designers consider different arrangements of lenses—i.e., topologies—and optimize the diameter, curvature, material, coating, etc. for each component.
2 Decomposition-Aware Distributional Optimization (DADO)
We begin by reminding the reader of the three primary steps that are iterated in a naive EDA, which is unaware of any decomposability (e.g.brookes20_eda_em). From there, we describe how infusing this algorithm with awareness of the known decomposition of $f(x)$ stands to make the EDA more statistically efficient, after which we formally introduce DADO.
Intuitively, one can think of an EDA as having a search distribution, $p_{\theta}(x)$, that acts as a spotlight on the design space, which is iteratively moved toward areas of the space with high $f(x)$ in expectation. Modern EDAs parameterize the search distribution with neural network generative models, such as a Variational Autoencoder (kingma13_vae; e.g.,  brookes19_cbas). In a decomposition-unaware EDA, after having initialized the search distribution, $p_{\theta^{0}}(x)$, with initial parameters, $\theta^{0}$, the EDA proceeds by iterating through these three steps for $n=1\dots N$ or until some convergence criteria has been met ( larranaga01_eda; brookes20_eda_em; Sec. A.2):
  1. 1.
    Draw $K$ samples from the current search distribution, $\{x^{k}\}_{k=1}^{K}\sim p_{\theta^{n}}(x)$
  2. 2.
    Score each sample with the objective function to get its weight, $w^{k}=f(x^{k})$.
  3. 3.
    Update the search distribution parameters with weighted maximum likelihood estimation (MLE), using the weighted samples, $\theta^{n+1}=\arg\max_{\theta}\;\mathbb{E}_{\{x^{k}\}}[w^{k}\log p_{\theta}(x^{k})$].
The weighted MLE problem in step 3 is typically solved with gradient descent; we will choose to use only a fixed number of gradient steps, which can be theoretically justified through the equivalence of EDAs to Expectation-Maximization (brookes20_eda_em).
Now that a regular EDA is fresh in our minds, we can consider what it might mean to infuse it with knowledge of how $f(x)$ is decomposed, and how this might prove useful. Before discussing the general case, here we first build intuition by considering a simpler example, one that is “fully”-decomposable rather than only soft-decomposable. That is, consider $f(x)=C_{1}(\hat{x}_{1})+C_{2}(\hat{x}_{2})$ where the two meta-variables $\hat{x}_{1}$ and $\hat{x}_{2}$ do not share any original variables, $x_{i}$, corresponding to a junction tree without any edges, only independent nodes. In this setting, we can break the EDA down into two independent EDAs, each with reduced dimensionality compared to the original EDA. Specifically, we can employ a fully-factorized search distribution (“fully” again denoting that no variable appears in more than one factor), $p_{\theta}(x)=p_{\theta_{1}}(\hat{x}_{1})p_{\theta_{2}}(\hat{x}_{2})$. We modify the three steps of the EDA as follows:
  1. 1.
    Instead of drawing samples from one search distribution over all design variables, we sample each meta-variable from its own search distribution separately, $\hat{x}_{1}^{k}\sim p_{\theta^{n}_{1}}(\hat{x}_{1})$ and $\hat{x}_{2}^{k}\sim p_{\theta^{n}_{2}}(\hat{x}_{2})$.
  2. 2.
    Instead of scoring $x^{k}$ with $f(x)$, we score each meta-variable separately with its component function, for $i\in\{1,2\}$, to obtain separate weights.
  3. 3.
    Instead of one update to a joint search distribution, we update each factor independently of the others with weighted MLE on samples from the relevant search distribution, $\theta_{i}^{n+1}=\arg\max_{\theta_{i}}\;\mathbb{E}_{\{\hat{x}_{i}^{k}\}}\left[w_{i}^{k}\log p_{\theta_{i}}(\hat{x}_{i}^{k})\right]$, for $i\in\{1,2\}$.
What has this bought us? We have split one optimization problem over a combinatorial space into two independent optimization problems, each over a smaller combinatorial space. For example, for binary design variables, where each of the two meta-variables comprises say 15 variables, we have transformed one joint optimization over a space of size $2^{30}\approx 10^{9}$ to two independent optimization problems each of size $2^{15}\approx 10^{5}$, thereby reducing the effective search space size by four orders of magnitude. For larger alphabets and sequence lengths, these differences will be larger still. Importantly, the weighted MLE update with a fixed number of samples, $K$, is more statistically efficient for each lower-dimensional search distribution factor than it is for a decomposition-unaware EDA using a single joint search distribution over all design variables.
In the general case, that is, beyond full decomposability, the same variable, $x_{i}$, can participate in multiple component functions $C_{j}(\hat{x}_{j})$. In other words, the meta-variables $\hat{x}_{j}$ may overlap, corresponding to there being edges in our junction tree. Because of this “component coupling”, the EDA can no longer independently optimize each component function and will require coordination between meta-variables. That is, we can no longer simply divide and conquer as above; rather, we must divide, and then collaborate to conquer. Collaboration among meta-variables will occur by way of a new form of message-passing that allows us to perform distributional optimization while leveraging the decomposability reflected by the junction tree.
Earlier we described decomposability as $f(x)=C_{1}(\hat{x}_{1})+C_{2}(\hat{x}_{2}),\dots,C_{\kappa}(\hat{x}_{\kappa})$ to help provide some intuition. However, to develop DADO it will be more convenient to equivalently describe the decomposability with a junction tree defined by nodes that contain multiple design variables and undirected edges that connect the nodes. We will define this junction tree decomposition in Sec. 2.1 and explain its connection to the above decomposability in Sec. A.1. It suffices to say here that design variables $x_{i}$ that occurred uniquely in one meta-variable $\hat{x}_{j}$ now sit in a single node, $\tilde{x}_{k}$; variables that previously occurred in multiple meta-variables now appear in multiple adjacent nodes. With the idea of a junction tree in hand, we are now set to complete our intuitive overview of DADO, which will use a factorized search distribution matching the topology of the junction tree. Namely, we will convert the undirected junction tree into a directed, acyclic graph (DAG) by rooting it and then directing the edges outward from the root. This operation will yield a search distribution that can be written as $p_{\theta}(x)=p_{\theta}(\tilde{x}_{r})\sum_{i}{p_{\theta}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})}$, for root node $\tilde{x}_{r}$ and non-root nodes, $\tilde{x}_{i}$, with parents, $\tilde{x}_{\text{p}(i)}$. Note that each non-root node has exactly one parent in a DAG. Each search distribution factor is a conditional probability distribution that will be parameterized by a neural network such as an autoregressive model.
Consequently, sampling from such a factorized search distribution is both computationally efficient and easy, so step 1 of the EDA remains straightforward. However, steps 2 and 3 of the EDA are not so easily generalized. For these steps to make use of the decomposability, we will need to leverage a special kind of message-passing to coordinate updates to different factors of the search distribution so as to achieve a globally consistent update for the entire search distribution. At a high level, our decomposition-aware EDA will proceed as follows for the $n+1^{\text{th}}$ iteration, compared to the fully-decomposable EDA presented previously:
  1. 1.
    Instead of sampling each meta-variable from an independent search distribution factor, we sample meta-variables sequentially from conditional search distribution factors, starting with the root node, $\tilde{x}_{r}^{k}\sim p_{\theta^{n}}(\tilde{x}_{r})$, and conditionally sampling down to the leaves, $\tilde{x}_{i}^{k}\sim p_{\theta^{n}}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)}^{k})$.
  2. 2.
    Instead of scoring each meta-variable $\tilde{x}_{i}^{k}$ separately with its component function, $C_{i}(\tilde{x}_{i}^{k})$, we use message-passing on the junction tree to obtain scores accounting for each meta-variable’s coupling to its parents and children, $w_{r}^{k}=Q_{r}(\tilde{x}_{r}^{k})$ for the root and $w_{i}^{k}=Q_{i}(\tilde{x}_{i}^{k},\tilde{x}_{\text{p}(i)}^{k})$ for non-root nodes.
  3. 3.
    We perform a coordinated update of the search distribution via per-factor weighted MLE on the relevant samples, $\theta_{r}^{n+1}=\arg\max_{\theta_{r}}\;\mathbb{E}_{\{\tilde{x}_{r}^{k}\}}\left[w_{r}^{k}\log p_{\theta}(\tilde{x}_{r}^{k})\right]$ for the root and $\theta_{i}^{n+1}=\arg\max_{\theta_{i}}\;\mathbb{E}_{\{\tilde{x}_{i}^{k},\tilde{x}_{\text{p}(i)}^{k}\}}\left[w_{i}^{k}\log p_{\theta}(\tilde{x}_{i}^{k}\mid\tilde{x}_{\text{p}(i)}^{k})\right]$ for non-root nodes. These updates can be performed separately, as in the fully-decomposed EDA, but are coupled according to the junction tree by the weights.
Having provided an intuitive overview of DADO (Fig. 1), we next provide a more formal exposition.
2.1 Formal exposition of DADO
The goal of DADO is to obtain a generative model, $p_{\theta}(x)$, that maximizes the EDA objective, $\arg\max_{\theta}\mathbb{E}_{p_{\theta}(x)}[f(x)]$, while leveraging decomposability in $f(x)$ for optimization efficiency. DADO requires as input a decomposed version of $f(x)$, which can be described by an undirected junction tree, $\mathcal{T}:=(\mathcal{N},\mathcal{E})$, with nodes, $\mathcal{N}$, and edges, $\mathcal{E}$. As noted in the introduction and shown in our experiments on proteins, identifying useful decomposability is feasible in practice. Given the junction tree decomposition, we write $f(x)=\sum_{i\in\mathcal{N}}f_{i}(\tilde{x}_{i})+\sum_{(i,j)\in\mathcal{E}}f_{i,j}(\tilde{x}_{i},\tilde{x}_{j})$, where we refer to $f_{i}(\tilde{x}_{i})$ as “node component functions” and $f_{i,j}(\tilde{x}_{i},\tilde{x}_{j})$ as “edge component functions” (Fig. 1a). This functional structure is intimately related to notions of the “epistatic landscape” of a protein’s property function and spectral analysis (see Sec. A.1.1). When the component functions are not known a priori, they can be parameterized and fit to labeled data.
We will begin our exposition by recalling how to do decomposition-aware exact (non-distributional) optimization, that is, to solve $\arg\max_{x}f(x)$. This problem is efficiently solved with a classical message-passing algorithm, which coordinates local optimizations across parts of the junction tree to obtain a single global optimum. Its efficiency comes from breaking optimization over all variables jointly into separate optimizations for each (smaller) meta-variable. Having loaded the reader with this intuition, we then adapt these ideas to distributional optimization, yielding DADO.
2.1.1 Classical message-passing for non-distributional optimization
Although classical message-passing has been largely used for probabilistic inference on probabilistic graphical models (pearl88_bp_textbook; shah_14_mit_inference-algorithms_message-passing), it can also be used for exact optimization of a function, $x^{*}=\arg\max\nolimits_{x}f(x)$ (vlassis04_maxplus). In particular, message-passing can be used to find a global maximum of a function defined on an undirected junction tree, $\mathcal{T}$, by making use of its topology for optimal time-complexity. In particular, one takes the junction tree, which is undirected, and roots it to obtain a directed tree, which induces a hierarchy among the meta-variables from root to leaves. Each node is responsible for accumulating information from all nodes in its sub-tree and then passing this information on to its parent. Consequently, the root node receives information from the entire tree, which is sufficient to set its variables in a globally optimal manner. Then, starting with the root, each parent communicates its variables’ optimal settings to its children, which can in turn set their variables optimally, and so forth.
To obtain the rooted tree, $\mathcal{T^{\prime}}:=(\mathcal{N^{\prime}},\mathcal{E^{\prime}})$, from the junction tree, one keeps the same nodes, $\mathcal{N^{\prime}}=\mathcal{N}$, chooses a root node, $r$, and directs all edges in $\mathcal{E}$ outward from $r$, yielding directed edges, $\mathcal{E^{\prime}}$. Although rooting at any node will suffice, we choose $r$ such that $\mathcal{T^{\prime}}$ has the shortest height possible. Message-passing finds a global optimum in two passes through the tree: one round of passing messages up from leaves to root, and then one round passing messages back down to the leaves.
Given $\mathcal{T^{\prime}}$, classical message-passing first uses dynamic programming to accumulate information from the leaves up to the root. Similarly to any dynamic programming procedure, we accumulate solutions to increasingly larger intermediate sub-problems. In this case, each sub-problem is to find the value of $f(x)$ evaluated on only a subset of meta-variables, rather than on the full set of variables in $x$. Each sub-problem is tractable owing to the decomposition of the objective function into component functions for each node and edge, and by respecting the partial order of sub-problems induced by $\mathcal{T^{\prime}}$. Specifically, one computes a value function, $V^{\text{max}}_{i}(\tilde{x}_{\text{p}(i)})$, for each node $i\in\mathcal{N^{\prime}}\setminus\{r\}$, which tells us for each setting of its parent, $\tilde{x}_{\text{p}(i)}$, the value of the intermediate objective function defined by the edge component function, $f_{\text{p}(i),i}(\tilde{x}_{\text{p}(i)},\tilde{x}_{i})$, plus all component functions over the sub-tree rooted at $i$, given that all nodes maximize their respective intermediate objectives. Computing value functions comprises the first pass through the tree, from leaves to root:
$V^{\text{max}}_{i}(\tilde{x}_{\text{p}(i)}):=\max\nolimits_{\tilde{x}_{i}}\Big(f_{i}(\tilde{x}_{i})+f_{\text{p}(i),i}(\tilde{x}_{\text{p}(i)},\tilde{x}_{i})+\sum\nolimits_{c\in\text{children}(i)}V^{\text{max}}_{c}(\tilde{x}_{i})\Big).$
Notably, $V^{\text{max}}_{i}(\tilde{x}_{\text{p}(i)})$ provides sufficient information about all nodes in the sub-tree rooted at $i$ to optimally choose the value of $\tilde{x}_{\text{p}(i)}$ with respect to its children. Thus it follows that once all value functions have been computed, the root’s assignment can be set in a globally optimal manner from its children’s value functions,
$\tilde{x}_{r}^{*}:=\arg\max\nolimits_{\tilde{x}_{r}}\Big(f_{r}(\tilde{x}_{r})+\sum\nolimits_{c\in\text{children}(r)}V_{c}(\tilde{x}_{r})\Big).$
Having chosen the root assignment, we then pass it down the tree as $\tilde{x}_{\text{p}(i)}=\tilde{x}_{\text{p}(i)}^{*}$ to its children, which successively pass their chosen assignments to their children, all the way to the leaves,
$\tilde{x}_{i}^{*}:=\arg\max\nolimits_{\tilde{x}_{i}}\Big(f_{i}(\tilde{x}_{i})+f_{\text{p}(i),i}(\tilde{x}_{\text{p}(i)}^{*},\tilde{x}_{i})+\sum\nolimits_{c\in\text{children}(i)}V_{c}(\tilde{x}_{i})\Big),$
resulting in a global maximizer $x^{*}$ of $f(x)$. This dynamic programming “traceback” of optimal assignments back down the tree constitutes our second and final pass of messages.
Alternative notation.
For convenience of our generalization to distributional optimization, we re-write the parent value functions $V^{\text{max}}_{i}(\tilde{x}_{\text{p}(i)})$ in terms of child-parent, $Q^{\text{max}}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})$, and single-node, $Q^{\text{max}}_{i}(\tilde{x}_{i})$ value functions:
$V^{\text{max}}_{i}(\tilde{x}_{\text{p}(i)})$ $:=\max\nolimits_{\tilde{x}_{i}}\;Q^{\text{max}}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)}),\text{\;\;where}$ (1)
$Q^{\text{max}}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})$ $:=f_{\text{p}(i),i}(\tilde{x}_{\text{p}(i)},\tilde{x}_{i})+Q^{\text{max}}_{i}(\tilde{x}_{i})\text{\;\;and}$ (2)
$Q^{\text{max}}_{i}(\tilde{x}_{i})$ $:=f_{i}(\tilde{x}_{i})+\sum\nolimits_{c\in\text{children}(i)}V^{\text{max}}_{c}(\tilde{x}_{i}).$ (3)
In contrast to the original parent value functions, $Q^{\text{max}}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})$ represents the effect of the choice of both $\tilde{x}_{\text{p}(i)}$ and $\tilde{x}_{i}$ on their edge component function plus all component functions over the sub-tree rooted at $i$, assuming all descendants of $i$ maximize their corresponding value functions. Intuitively, $Q^{\text{max}}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})$ is the value function on edge $(\text{p}(i),i)$ prior to $\tilde{x}_{i}$ being maximized out, which will become useful if we want to, say, sample $\tilde{x}_{i}$ according to some distribution instead. $Q^{\text{max}}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})$ is composed of two terms, one of which depends on its parent, and one of which doesn’t, $Q^{\text{max}}_{i}(\tilde{x}_{i})$. Written using the $Q$-functions just defined, the equivalent traceback equations for selecting a globally optimal assignment, $x^{*}$, are simply $\tilde{x}_{r}^{*}:=\arg\max_{\tilde{x}_{r}}Q^{\text{max}}_{r}(\tilde{x}_{r})$ and $\tilde{x}_{i}^{*}:=\arg\max_{\tilde{x}_{i}}Q^{\text{max}}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)}^{*})$. In other words,
$x^{*}=\arg\max\nolimits_{x}f(x)$ $=\{\arg\max\nolimits_{\tilde{x}_{r}}Q^{\text{max}}_{r}(\tilde{x}_{r})\}\;+\;\{\arg\max\nolimits_{\tilde{x}_{i}}Q^{\text{max}}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)}^{*})\}_{(\text{p}(i),i)\in\mathcal{E^{\prime}}}$ (4)
$=\arg\max\nolimits_{x}\Big(Q^{\text{max}}_{r}(\tilde{x}_{r})+\sum\nolimits_{(\text{p}(i),i)\in\mathcal{E^{\prime}}}Q^{\text{max}}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})\Big).$ (5)
2.1.2 From classical message passing to distributional optimization
In the same way that an EDA transforms $\arg\max_{x}f(x)$ into a distributional optimization problem, we can rewrite the equivalent message-passing objective in Eq. 5 as DO Because the original optimization problems over $x$ are equivalent, their DO formulations are equivalent too,
$\arg\max\nolimits_{\theta}\mathbb{E}_{p_{\theta}(x)}[f(x)]=\arg\max\nolimits_{\theta}\Big(\mathbb{E}_{p_{\theta}(x)}[Q^{\text{max}}_{r}(\tilde{x}_{r})]+\sum\nolimits_{(\text{p}(i),i)\in\mathcal{E^{\prime}}}\mathbb{E}_{p_{\theta}(x)}[Q^{\text{max}}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})]\Big),$ (6)
where we’ve used linearity of expectations on the right side. However, a generic joint search distribution cannot take advantage of the linear additivity in value functions over the junction tree topology. That is, while the classical traceback equations perform maximization over each meta-variable separately, Eq. 6 uses a single, unfactorized search distribution over all variables, $p_{\theta}(x)$, to optimize each $Q$-function. We address this next, by factorizing the search distribution.
Factorized search distribution.
Classical message-passing (Eq. 4) independently maximizes each $Q$-function conditional on the choice of $\tilde{x}_{\text{p}(i)}$, instead of explicitly maximizing all variables in $x$ jointly (Eq. 5). This is possible because each $Q_{i}$ captures all relevant global information needed to choose $\tilde{x}_{i}$. It stands to reason then that DO can do something similar. Specifically, instead of training a single joint search distribution, we train smaller search distributions over each $\tilde{x}_{i}$, conditional on $\tilde{x}_{\text{p}(i)}$, to separately maximize each $Q^{\text{max}}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})$. That is, we factor the search distribution according to $\mathcal{T^{\prime}}$, resulting in a DAG search distribution, $p_{\theta}(x)\coloneqq p_{\theta}(\tilde{x}_{r})\prod_{(\text{p}(i),i)\in\mathcal{E^{\prime}}}p_{\theta}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})$, for root node $r$, non-root nodes $i$, and parents $\text{p}(i)$, connected by directed edges $\mathcal{E^{\prime}}$, and with parameters, $\theta$. This factorized search distribution can be plugged into Eq. 6 for an equivalent optimization problem. We refer to each element of this product as one of the factors of the search distribution. In our implementation, each factor has completely separate parameters, though we write a shared $\theta$ for conciseness. Notice that each factor of the search distribution interacts with each other factor through the directed edges, $\mathcal{E^{\prime}}$. That is, the distribution of $\tilde{x}_{i}$ depends on its parent’s factor, and through it, all of its parent’s ancestors: $p_{\theta}(\tilde{x}_{i})=p_{\theta}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})p_{\theta}(\tilde{x}_{\text{p}(i)})$. In turn, each of node $i$’s children’s factors depends on $p_{\theta}(\tilde{x}_{i})$. Due to this coupling, we cannot optimize each factor fully independently. But we can still update each factor separately from the others in a globally-consistent manner via message-passing. In particular, each factor will only be responsible for directly optimizing its own meta-variable, but will need to coordinate with its neighboring factors by getting information from them about how they are optimizing their meta-variables in a manner analogous to classical message-passing. Our current messages, $Q_{i}^{\text{max}}$, convey the value of each intermediate objective when all meta-variables are chosen via maximization. For DO, we’ll require messages that communicate values when meta-variables are chosen according to the factorized search distribution, $p_{\theta}(x)$.
Distributional value functions.
While one certainly could choose to optimize classical value functions using an EDA search distribution (Eq. 6), it doesn’t make sense for two main reasons. As we just mentioned, DO aims to train $p_{\theta}(x)$ such that it maximizes $f(x)$, or equivalently, the sum of $Q$-functions, in expectation. Therefore, the DO objective should consider intermediate objective values for meta-variables chosen according to the current search distribution, not those chosen by explicit maximization, as in $V^{\text{max}}_{i}(\tilde{x}_{\text{p}(i)})$. Additionally, computing classical value functions requires enumerating all assignments of each $\tilde{x}_{i}$ for the maximum used in $V^{\text{max}}_{i}(\tilde{x}_{\text{p}(i)})$. When $\tilde{x}_{i}$ contains more than a few design variables, this max operation quickly becomes intractable. One reason for doing distributional optimization is to avoid enumerating massive design spaces. To address both issues, we define corresponding distributional value functions that fulfill both desiderata:
$V^{\theta}_{i}(\tilde{x}_{\text{p}(i)})$ $:=\mathbb{E}_{p_{\theta}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})}[Q^{\theta}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})],\text{\;\;where}$ (7)
$Q^{\theta}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})$ $:=f_{\text{p}(i),i}(\tilde{x}_{\text{p}(i)},\tilde{x}_{i})+Q^{\theta}_{i}(\tilde{x}_{i})\text{\;\;and}$ (8)
$Q^{\theta}_{i}(\tilde{x}_{i})$ $:=f_{i}(\tilde{x}_{i})+\sum\nolimits_{c\in\text{children}(i)}V^{\theta}_{c}(\tilde{x}_{i}).$ (9)
Compared to Eq. 1, the distributional $V$-functions compute the value in expectation under $p_{\theta}$ instead of a $\max$. Because they use an expectation instead of a $\max$ operation, these value functions can be approximated tractably and without bias by drawing Monte-Carlo samples from the search distribution. Moreover, each distributional value function lower bounds each corresponding classical value function because the expectation of a function cannot exceed its maximum (details in Sec. A.3). As a result, the sum of classical value functions in the objective in Eq. 6 is bounded below by the sum of distributional value functions, which we’ll optimize instead:
$\mathbb{E}_{p_{\theta}(x)}[Q^{\text{max}}_{r}(\tilde{x}_{r})]+\sum_{(\text{p}(i),i)\in\mathcal{E^{\prime}}}\mathbb{E}_{p_{\theta}(x)}[Q^{\text{max}}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})]\geq\mathbb{E}_{p_{\theta}(x)}[Q^{\theta}_{r}(\tilde{x}_{r})]+\sum_{(\text{p}(i),i)\in\mathcal{E^{\prime}}}\mathbb{E}_{p_{\theta}(x)}[Q^{\theta}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})].$ (10)
Distributional optimization with value functions.
All that remains is to derive an update rule that treats each search distribution factor separately. Since each expectand in Eq. 10 doesn’t depend on descendant meta-variables, we can replace $p_{\theta}(x)$ with each meta-variable’s marginal distribution:
$\arg\max_{\theta}\mathbb{E}_{p_{\theta}(\tilde{x}_{r})}[Q^{\theta}_{r}(\tilde{x}_{r})]+\sum\nolimits_{(\text{p}(i),i)\in\mathcal{E^{\prime}}}\mathbb{E}_{p_{\theta}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})p_{\theta}(\tilde{x}_{\text{p}(i)})}[Q^{\theta}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})].$ (11)
We then follow the EDA derivation to arrive at an update rule (details in Sec. A.3) in which each term is optimized by only a single factor; dependence on $p_{\theta}(\tilde{x}_{\text{p}(i)})$ is approximated by sampling. We write DADO’s update rule (sharing a single set of samples; see Fig. 1b) as a sum of weighted likelihoods for each search distribution factor,
$\theta^{n+1}=\arg\max_{\theta}\sum\nolimits_{x\sim p_{\theta^{n}}(x)}\Big(Q_{r}^{\theta^{n}}(\tilde{x}_{r})\log p_{\theta}(\tilde{x}_{r})+\sum\nolimits_{(\text{p}(i),i)\in\mathcal{E^{\prime}}}Q_{i}^{\theta^{n}}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})\log p_{\theta}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})\Big),$ (12)
which can be written equivalently as separate updates because the factors don’t share parameters, as desired:
$\theta_{r}^{n+1}$ $=\arg\max_{\theta_{r}}\sum\nolimits_{x\sim p_{\theta^{n}}(x)}Q_{r}^{\theta^{n}}(\tilde{x}_{r})\log p_{\theta_{r}}(\tilde{x}_{r})\text{, \;\; and}$ (13)
$\theta_{i}^{n+1}$ $=\arg\max_{\theta_{i}}\sum\nolimits_{x\sim p_{\theta^{n}}(x)}Q_{i}^{\theta^{n}}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})\log p_{\theta_{i}}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)}),\;\;\forall\;i\in\mathcal{N}\setminus\{r\}.$ (14)
Each factor is weighted by its corresponding value function, enabling it to coordinate with all its descendant factors despite their being updated separately. The whole DO is tied together at the top by the root factor. We emphasize that DADO’s update is more statistically efficient than a naive EDA’s because DADO gets to use all $K$ samples for weighted MLE on each lower-dimensional factor distribution (i.e., ratio of number of samples to number of dimensions is larger). Our resulting algorithm (Alg. 1) fits into the three EDA steps: (1) designs are sampled from $p_{\theta^{n}}(x)$, (2) weights, here each $Q$-function instead of $f(x)$, are computed, and (3), each search distribution factor receives its own independent weighted maximum likelihood update based on these weights (steps 1 and 2 illustrated in Fig. 1b). These are repeated until convergence, or for some fixed number of iterations. The factor updates are coupled only through the $Q$-functions, the messages across edges. $\{Q^{\theta^{n}}_{i}\}_{i\in\mathcal{N}}$ are only valid while $\theta$ is close to $\theta^{n}$, meaning one must balance how many gradient steps are taken before drawing new samples. If too many gradient steps are taken without resampling, a factor may be changing its parameters to collaborate with another factor which has already changed its behavior. It’s common for EDAs to include an additional hyperparameter $W(\cdot)$, a monotonic shaping function applied to the weights, to alter optimization dynamics (brookes20_eda_em), so we include it in Alg. 1 on lines 16 and 18. Choosing $W$ to be the identity recovers our derivation above. Notice that in place of the classical message-passing traceback equations, DADO simply performs sequential conditional sampling from its search distribution. Interestingly, a very similar algorithm can be derived without using a lower bound on the value functions (Eq. 10) if one adds an entropy-maximizing term to the initial DO objective (Sec. A.4).
Algorithm 1 Decomposition-Aware Distributional Optimization (DADO)
1:Given: Decomposition $\left(\text{junction tree}\;\mathcal{T}=(\mathcal{N},\mathcal{E}),\;\text{component functions}\;\{f_{i,j}(\tilde{x}_{i},\tilde{x}_{j})\}_{(i,j)\in\mathcal{E}}\;\text{and}\;\{f_{i}(\tilde{x}_{i})\}_{i\in\mathcal{N}}\right)$
2:Root $\mathcal{T}$ such that tree height is minimized, yielding directed junction tree, $\mathcal{T^{\prime}}=(\mathcal{N},\mathcal{E^{\prime}})$
3:Factorize search distribution according to $\mathcal{T^{\prime}}$ as $p_{\theta}(x)\coloneqq p_{\theta}(\tilde{x}_{r})\prod_{(\text{p}(i),i)\in\mathcal{E^{\prime}}}p_{\theta}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})$; initialize $\theta$
4:Sort nodes topologically: order $O_{\mathcal{N}}$ such that $O_{\mathcal{N}}^{(0)}$ is a leaf index and $O_{\mathcal{N}}^{(L-1)}$ is root index
5:for iteration $n=1,2,\ldots,N$ do
6:  $\{\tilde{x}^{k}\}_{k=1}^{K}\sim p_{\theta}(x)\triangleright$ Sample from the factorized search distribution
7:  for $i\in O_{\mathcal{N}}$ do $\triangleright$ Estimate distributional value functions from samples
8:   for $k=1,2,\ldots K$ do
9:     $Q_{i}(\tilde{x}_{i}^{k})\leftarrow f_{i}(\tilde{x}_{i}^{k})+\sum_{c\in\text{children}(i)}\sum_{k=1}^{K}[Q_{c}(\tilde{x}_{c}^{k},\tilde{x}_{i}^{k})]$
10:     if $i$’s parent, p(i)(i), exists:
11:      $Q_{i}(\tilde{x}_{i}^{k},\tilde{x}_{\text{p}(i)}^{k})\leftarrow f_{\text{p}(i),i}(\tilde{x}_{\text{p}(i)}^{k},\tilde{x}_{i}^{k})+Q_{i}(\tilde{x}_{i}^{k})$
12:   end for
13:  end for
14:  for all $i\in\mathcal{N}$, in parallel do $\triangleright$ Update each search distribution factor using value functions as weights
15:   if $i=r$:
16:    $\theta_{r}\leftarrow\arg\max\nolimits_{\theta_{r}}\sum\nolimits_{k=1}^{K}W(Q_{r}(\tilde{x}_{r}^{k}))\log p_{\theta_{r}}(\tilde{x}_{r}^{k})$
17:   else
18:    $\theta_{i}\leftarrow\arg\max\nolimits_{\theta_{i}}\sum\nolimits_{k=1}^{K}W(Q_{i}(\tilde{x}_{i}^{k},\tilde{x}_{\text{p}(i)}^{k}))\log p_{\theta_{i}}(\tilde{x}_{i}^{k}\mid\tilde{x}_{\text{p}(i)}^{k})$
19:  end for
20:end for
3 Related work
The closest work to ours in the EDA community is the Factorized Distribution Algorithm (FDA), wherein a decomposition is leveraged to factorize the conditional probability table (CPT) that parameterizes the search distribution (muhlenbein99_fda). However, FDA cannot leverage message-passing to give different weights to different variables so as to more efficiently update the search distribution. In related work, pelikan05_boa replaced the CPTs with Bayesian networks, but still had no means to coordinate among variables through message-passing or the like.
In a complementary line of work on policy optimization in reinforcement learning (analogous to learning the search distribution in EDAs), coordination between variables is enabled by message-passing on a factorized search distribution/policy, but this line of work is only suitable for graph topologies that are Markov chains—that is, chain graphs. These graphs arise from reward functions, $f(x)$, that decompose in time according to a first-order Markov assumption (peters07_rwrpeng19_awrnair20_awac). DADO can be viewed as generalizing policy optimization over chains to arbitrary graphs. We explore this connection in more detail in Sec. A.4.
Adjacent to our problem of interest, because they focus solely on real-valued design spaces,  grudzien24_fgm introduces a new way to discover a function decomposition from data and then demonstrates that using it improves out-of-distribution generalization and thus, downstream optimization performance. Separately, the Bayesian optimization (BO) community has developed methods for dynamically inferring a function decomposition from data during active learning. Although they make use of classical message-passing for coordination, they do not employ distributional optimization, nor can these methods be readily generalized to do so (kandasamy15_additive-BOhan21_tree-BOhoang18_JT-BOrolland18_JT-BObardou24_loopy_BO). This community has shown that using approximate or even random decompositions can be helpful (ziomek23_random-decompositions_BO). DADO could, in principle, be used within the BO inner loop, although such an investigation is beyond the scope of the present work.
4 Experimental results
We perform two sets of experiments with increasing resemblance to real-world scientific design. In each setting we compare a naive EDA, which is unaware of the function decomposition, to DADO, which is aware of it. In the first setting, we create synthetic functions $f(x)$, while in the second setting, we focus on functions derived from real protein data. In all experiments we designed sequences of fixed length $L$, where each position is one of $D=20$ amino acids. We used $N=100$ EDA training iterations and $G=1$ gradient steps for each weighted maximum likelihood update. For the search distribution of DADO, we used MLP-based autoregressive neural networks for each node—specifically, each variable $x_{j}$ in $\tilde{x}_{i}$ is modeled by an MLP of size $[64,64]$ that takes as input all conditioning variables (i.e., $\tilde{x}_{\text{p}(i)}$ and any preceding variables in $\tilde{x}_{i}$) and outputs $D$ logits. Within $\tilde{x}_{i}$, an arbitrary decoding order was used, and any overlapping variables, that is, $x_{j}$ in both $\tilde{x}_{i}$ and $\tilde{x}_{\text{p}(i)}$, were considered the purview of the parent and thus omitted from the child except as conditioning. Our naive EDA implementation is the same, only assuming all variables are in one meta-variable—hence, a standard autoregressive model over all variables. For gradient descent to fit the search distribution, we used the AdamW optimizer (loshchilov17_adamw) with default momentum parameters $\beta_{1}=0.9,\beta_{2}=0.999$. All code is implemented using Jax, Flax, and Optax (deepmind20_jax). For the shaping function we choose ${W}(s)=\exp\frac{s}{\beta}$ and tuned the temperature, $\beta$. To choose $\beta$ and the learning rate, $\eta$, we performed a hyperparameter sweep separately for each method and each $f(x)$, taking $(\eta^{*},\beta^{*})$ with the largest sample mean at the last iteration. For synthetic experiments, we swept 100 combinations between $\eta\in[10^{-5},\;5\times 10{-2}]$ and $\beta\in[0.1,\;8]$. For protein experiments, we swept only 54 pairs between $\eta\in[5\times 10^{-5},\;5\times 10{-2}]$ and $\beta\in[0.01,\;5]$, for expediency. We confirmed that neither method’s $\eta^{*}$ or $\beta^{*}$ was on a boundary of the sweep. $(\eta^{*},\beta^{*})$ were then used for all replicate runs over random seeds dictating the initial search distribution and sampling. Although not essential to our method as described in Alg. 1, in our implementation we reduce the variance of the search distribution update by using mean-shifted $Q$-functions, $Q^{\theta}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})-\mathbb{E}_{\{\tilde{x}_{i}^{k}\}}[Q^{\theta}_{i}(\tilde{x}_{i}^{k},\tilde{x}_{\text{p}(i)})]$, which are unbiased (williams92_reinforce). Statistical significance of differences in performance is computed by comparing the area under the curve (AUC) of $f(x)$-sample mean over iterations curves, using paired two-sided t-tests over random seeds.
4.1 Comparison on synthetic functions
[…]
Figure 2: Comparison of a naive EDA to DADO on synthetic problems. We created three random functions, $f(x)$, each with a randomly chosen junction tree decomposition with maximum node size of one, and randomly chosen parameters. Each experiment used alphabet size $D=20$ and sequence length, a, $L=25$, b, $L=50$, and c, $L=200$. Each of the two methods drew $K=100$ samples per iteration. For each iteration, we show the mean (solid line) and 95% confidence interval (shaded envelope) of the 100 samples evaluated on $f(x)$, averaged across results from 20 random seeds. P-values are from two-sided paired t-tests that AUC of the per-iteration mean is different between methods, using the 20 mean curves.
Here we used junction trees with meta-variable nodes containing only one original variable, i.e., an exact tree decomposition. We did so because this allowed us to better control the difficulty of the synthetic functions in that we could specify each edge component function as a table of size $D^{2}$. That is, an edge component function in which each node has only a single variable and thus can take on $D$ states each, is fully specified by a value for every possible pair of node choices, i.e., a table. For larger meta-variables, we would have required a more compact representation than a table (which scales exponentially with the number of variables in each meta-variable) such as a neural network-based function. However, randomly initialized neural networks tend to be overly smooth and unrealistically easy to optimize, and neural network outputs cannot generally be set in a controlled manner across the entire input domain due to its combinatorial size, making it nontrivial to obtain realistically difficult functions. We simulated one $f(x)$ for each of three sequence lengths, $L=\{25,50,200\}$, by first randomly sampling a junction tree topology, and then randomly specifying the component functions. Each node function, a $D$-vector, was sampled $f_{i}\in\mathbb{R}^{D}\sim\mathcal{N}(0,0.01)$, where $\mathcal{N}$ denotes a Gaussian distribution. Each edge function, a $D\times D$ matrix, was sampled, $f_{i,j}\in\mathbb{R}^{D\times D}\sim\mathcal{N}(0,0.0025)$. To ensure a reasonable degree of non-smoothness in $f(x)$, we further explicitly added what in biology is known as reciprocal sign epistasis (starr16_epistasis_evolution; li24_mlde_epistasis). Specifically, for each edge function we, twice, randomly assigned one of the 20 alphabet letters to each node, i) $x_{i}:=A,x_{j}:=B$ and ii) $x_{i}:=C,x_{j}:=D$. Next we sampled an effect size, $\lambda\sim\mathcal{N}(0,4)$. Finally, we set $f_{i,j}(x_{i}=A,x_{j}=B)=0$ and $f_{i,j}(x_{i}=C,x_{j}=B)=\lambda$, and also half of the time, $f_{i,j}(x_{i}=A,x_{j}=D)=0$ and $f_{i,j}(x_{i}=A,x_{j}=D)=\lambda$. Initial search distribution weights were drawn from $\mathcal{N}(0,0.0004)$ and the biases set to 0.
Both methods draw $K=100$ samples from their search distribution at each iteration. For all three sequence lengths, DADO outperforms the naive EDA, both visually, and in terms of a statistical test (Fig. 2). The final iteration is more concentrated in regions of high $f(x)$, and DADO also reaches these regions more quickly than the naive EDA, which we quantify with a statistical test on the area under the curve of the sample means. The difference is more pronounced for larger $L$, corresponding to larger design spaces. When we increase the number of samples drawn at each iteration to $K=1000$, both methods perform even better (Sec. A.6.1; Fig. A2).
4.2 Comparison on protein property predictors
[…]
Figure 3: Comparison of a naive EDA to DADO on protein property predictive models. For each of four proteins of varying length, a, Amyloid, b, AAV, c, GB1, and d, TDP-43, we fit a neural network property function, $f(x)$, adhering to a junction tree decomposition derived from the protein’s 3D structure, and then used a naive EDA and DADO to optimize them. Each approach drew $K=1000$ samples per EDA iteration. For each iteration, we show the mean (solid line) and 95% confidence interval (shaded envelope) of the 1000 samples evaluated on $f(x)$, averaged across results from 20 random seeds. P-values are from two-sided paired t-tests that AUC of the per-iteration mean is different between methods, using the 20 mean curves.
Here we anchor our experiments on $f(x)$ fit to real protein property datasets. We use four datasets describing aggregation (Amyloid), virus viability (AAV), binding affinity (GB1), and cell toxicity (TDP-43) details in App. A.5). For each dataset, we used AlphaFold3-predicted structures (abramson24_af3) on the wild-type sequence to obtain a 3D structure, from which we constructed a contact graph by thresholding the distance between pairs of residues with threshold $t$. Following brookes22_sparsity; romero13_GP_protein; voigt02_recombination, we use a threshold of $t=4.5\text{\AA }$. We interpret this contact map as a graph adjacency matrix, from which we algorithmically construct a junction tree (lauritzen88_JT). This defines the decomposition graph needed for DADO, but we must also fit the corresponding component functions on the protein assay-labeled data. To do so, we employ an MLP-based predictive model that strictly enforces the decomposition defined by the junction tree. In particular, we use an MLP that takes as input the sequence, and, critically, also a bit-vector specifying active variables. One function evaluation requires calling this MLP for every node and edge function and summing. We used 5-fold cross validation to sweep through hidden layers of size ($[16,16]$, $[128,16]$, or $[128,128,16]$), learning rate ($0.001$ or $0.0001$), and number of training iterations ($5,000$ or $50,000$), to choose hyperparameters with the lowest cross-validation mean-squared error. Finally, we trained the model with those hyper-parameters using all data available. The result, a predictive model, $f(x)$, can be evaluated on any sequence $x\in\mathcal{X}$. We initialized the search distribution by training on the 1,000 sequences in the protein property dataset with the lowest assay labels, for 1,000 iterations, with a learning rate of $2\times 10^{-3}$.
Both methods draw $K=1000$ samples from their search distribution at each iteration. For three out of four datasets (Amyloid, AAV, and GB1), DADO reaches areas of high $f(x)$ much more quickly than the decomposition-unaware EDA (Fig. 3). In the fourth case, TDP-43, both methods converge rapidly and there is no noticeable difference in performance, despite the protein being longer than the others and hence the search space combinatorially larger. This can happen if $f$ has many local optima with near-optimal values, making it very easy to optimize. Our experiments on a few more proteins also support the conclusion that DADO either outperforms the naive EDA or is comparable (Fig. A3). When we decrease $K$ to $100$ (Fig. A4), we observe an improvement over the EDA for some saturated problems (both methods converged to roughly the same $f(x)$ at $K=1000$), including TDP-43. The choice of $K$ can substantially affect how well a DO algorithm optimizes $f$; we find that using a larger $K$ tends to benefit both methods and that DADO performs better than or comparably to the decomposition-unaware EDA for all $K$ tested (Sec. A.6.2). Beyond $K$, when a practitioner chooses a decomposition for their predictive model, there will likely be choices that are more or less decomposed with similar predictive accuracy. All other things being equal, we expect that a more decomposed model will be easier to optimize, especially for a decomposition-aware method. In a case study on GB1, we test different choices of the contact threshold, $t$, and observe that a more decomposed model (smaller $t$) can be optimized much more quickly and effectively by DADO for a relatively small decrease in accuracy. (Sec. A.6.3; Fig. A5).
5 Discussion
We have proposed a new method for distributional optimization that can leverage arbitrary decomposability of the function being optimized. We have shown that DADO works as expected on synthetic problems–namely, better than a naive, decomposition-unaware EDA. We have also demonstrated DADO’s potential for practical utility on the problem of protein design. Importantly, it is not necessary that a problem strictly adhere to the specified decomposition in order to be useful. For instance, we found that by adjusting the contact threshold applied to AlphaFold3-predicted 3D structures, we could obtain a range of decomposed functions, including some with increased decomposability but comparable predictive accuracy. This result adds to observations in the BO community suggesting that approximate or even random decompositions can be helpful for optimization (e.g., kandasamy15_additive-BO; ziomek23_random-decompositions_BO). Obtaining similarly useful decomposability in other domains will require further investigation; however, as the real world typically is structured, we expect many areas will be amenable to doing so. Estimating decomposability from labeled data is an active area of research (poelwijk16_wht; grudzien24_fgm; park24_rfa).
It will be insightful to conduct future work interrogating the use of DADO with generalizations of the naive EDA, such as conditioning an unconditional generative model with an inaccurate property predictor in order to perform AI-guided design while accounting for the fact that $f(x)$ is not truly known everywhere due to finite data (brookes19_cbas). This particular problem may require distilling an unconditional model into one that respects the decomposition, although alternative strategies could prove useful. Additionally, DADO, like naive EDAs, is in principle agnostic to whether the design space is discrete or continuous. Investigating decomposability for continuous scientific design problems, such as those in trabucco22_design-bench, is an interesting direction for future work. Exploring extensions of policy optimization from RL (sutton98_rl-book) in the context of DADO could also be fruitful, such as using learned value functions. Finally, use of DADO within BO for optimization of the acquisition function is an exciting direction.
6 Reproducibility statement
We specify our experimental setup, hyperparameters, model architectures, and algorithm details in Sec. 4. Datasets used and their sources are detailed in App. A.5. Code for our method and experiments is available publicly on GitHub: https://github.com/james-bowden/DADO.
Acknowledgments
We thank Kuba Grudzien and members of the Listgarten group for helpful discussions regarding function decomposition, Hanlun Jiang and Peter Ma for discussions regarding scientific applications, and Toby Kreiman, Bear Xiong, and Aileen Zhang for feedback on the manuscript.
JCB was supported in part by an NSF GRFP fellowship and a UC Berkeley Chancellor’s fellowship. This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE 2146752. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. JCB and JL were supported in part by the Office of Naval Research (ONR) under grant N00014-23-1-2587, and the U.S. Department of Energy, Office of Science, Office of Biological and Environmental Research, Lawrence Livermore National Laboratory BioSecure SFA within the Secure Biosystems Design program (SCW1710).
Appendix A Appendix
A.1 Different ways of writing a function decomposition
In the introduction (Sec. 1), we define a decomposable function as one that can be written like $f(x)=C_{1}(\hat{x}_{1})+C_{2}(\hat{x}_{2}),\dots,C_{\kappa}(\hat{x}_{\kappa})$, with meta-variables $\hat{x}_{i}$ that are generic sets of original variables $x_{j}$. This formulation is most helpful for building intuition about fully-decomposable functions, in which there are no variables $x_{j}$ that appear in multiple meta-variables (i.e., meta-variables don’t overlap). It also encompasses functions for which there are overlapping meta-variables in the decomposition too.
$f(x)=C_{1}(\hat{x}_{1})+C_{2}(\hat{x}_{2}),\dots,C_{\kappa}(\hat{x}_{\kappa})$also provides the helpful intuition that someone looking to use DADO need only specify a decomposition at this level of detail—sets of design variables that directly interact—rather than as a graph. In the protein binding example given in Sec. 1, a user might specify three component functions: one over sequence positions in the binding interface, $\hat{x}_{0}$, one over sequence positions in the scaffold, $\hat{x}_{1}$, and one over a subset of positions tying together the binding interface and the scaffold, $\hat{x}_{2}$. Specifying a decomposition at this level may be easier than in its graph form. In the event that one has an exact decomposition given, this information can be represented either this way, or in terms of the graph introduced below. Situations in which an exact decomposition may be given include hardware design (e.g., due to manufacturing constraints, a topology of wires in a circuit is pre-specified) and bi-level optimizations (e.g., to design the best telescope, a scientist optimizes both the arrangement of lenses and their physical properties; an outer loop optimizes topologies and an inner loop using DADO evaluates a single topology by optimizing each lens’ parameters).
However, $f(x)=C_{1}(\hat{x}_{1})+C_{2}(\hat{x}_{2}),\dots,C_{\kappa}(\hat{x}_{\kappa})$ as written has no interpretation in terms of a graph, which is necessary for message-passing and for factorizing the search distribution (Sec. 2.1). To bridge this gap, we introduce a graph in which nodes are design variables and edges specify which design variables directly interact to influence $f$. Users of DADO may also specify the decomposition at this level; we expect this to be helpful when domain knowledge is available in the negative form, i.e., when the user can begin with the fully-connected graph and remove edges between design variables they think don’t directly interact. This graph can then be automatically transformed into a junction tree [lauritzen88_JT], which is what our algorithm actually takes as input. That is, DADO takes an undirected junction tree, $\mathcal{T}:=(\mathcal{N},\mathcal{E})$, with nodes, $\mathcal{N}$, and edges, $\mathcal{E}$. Each node is a set of design variable indices, such that indexing $x$ with this set yields a new meta-variable, $\tilde{x}_{i}$, correspondingly exactly to a node in the junction tree, whereas the old meta-variables $\hat{x}_{i}$ had no such interpretation. $f(x)$ must be decomposed instead according to $\mathcal{T}$.
We can rewrite the generic function decomposition, $f(x)=C_{1}(\hat{x}_{1})+C_{2}(\hat{x}_{2}),\dots,C_{\kappa}(\hat{x}_{\kappa})$, equivalently in terms of $\mathcal{N}$ and $\mathcal{E}$: $f(x)=\sum_{i\in\mathcal{N}}f_{i}(\tilde{x}_{i})+\sum_{(i,j)\in\mathcal{E}}f_{i,j}(\tilde{x}_{i},\tilde{x}_{j})$ (Fig. 1a). We’ve replaced component functions $C_{i}(\hat{x}_{i})$ on generic sets of design variable indices with either an equivalent “node component function”, $f_{i}(\tilde{x}_{i})$, or “edge component function”, $f_{i,j}(\tilde{x}_{i},\tilde{x}_{j})$. Correspondingly, $\hat{x}_{i}$ in the original formulation either corresponds exactly to some node in the junction tree, such that $\hat{x}_{i}=\tilde{x}_{j}$, or it corresponds to some edge in the junction tree, such that $\hat{x}_{i}=\tilde{x}_{j\;\cup\;k}$. The fully-decomposed, no-overlap EDA described in Sec. 2 corresponds to an edgeless junction tree—each component function is simply a node component function, $C_{i}(\hat{x}_{i})=f_{i}(\tilde{x}_{i})$, and there are no edge component functions—which is why a fully-factorized search distribution suffices. But if, for example, $\hat{x}_{0}=\{x_{0},x_{1}\}$ and $\hat{x}_{1}=\{x_{1},x_{2}\}$ overlap, then they must instead correspond to edge component functions. For a simplified case where the junction tree nodes have cardinality 1 (e.g., $\tilde{x}_{i}=x_{i}$), we might have $C_{0}(\hat{x}_{0})=f_{0,1}(\tilde{x}_{0},\tilde{x}_{1})$ and $C_{1}(\hat{x}_{1})=f_{1,2}(\tilde{x}_{1},\tilde{x}_{2})$. Fig. 1a only depicts edge component functions for clarity and simplicity; notice also that node component functions can technically be subsumed into appropriate edge component functions without loss of generality. The presence of edge component functions requires the coupling of the search distribution’s factors in a manner compatible with $\mathcal{E}$; a fully-factorized search distribution with independent distributions for each meta-variable will no longer suffice.
A.1.1 Connection to epistasis
Function decomposition corresponds intimately to notions of the epistatic landscape for a protein property function (e.g.poelwijk16_whtwu82_pottsotwinowski13_truncated-epistatic-regressionlipsh24_epistasis-protein-design). An “epistatic expansion” of $f$,
$f(x)=\beta+\sum_{i=0}^{L}\beta_{i}[x_{i}]+\sum_{i=0}^{L}\sum_{j>i}^{L}\beta_{i,j}[x_{i},x_{j}]+...+\beta_{0,1,...L-1}[x_{0},x_{1},...,x_{L-1}],$ (15)
is a decomposition of $f$ into a bias term, first-order terms, second-order terms, and so forth up to $L$-order epistasis. For the first-order terms, each $\beta_{i}$ is a vector which is indexed by the particular value of $x_{i}$. For higher-order terms, $\beta$ is a $d$-dimensional tensor. One can obtain a decomposition graph from an epistatic expansion as follows: for each nonzero term in the expansion, add an edge between each design variable in the term. A function in which only a few, lower-order terms were nonzero would be relatively easy to optimize.
An expansion of this sort can be thought of as a spectral decomposition of $f$ into components of lower and higher frequencies, and can be computed for small landscapes using a discrete Fourier transform. A function with primarily lower-frequency components would be smooth and easy to optimize whereas a function with higher-frequency components would be rugged and difficult to find the global optimum of. One can imagine a worst-case scenario, that of $L$-order epistasis, of which a needle-in-a-haystack function is an example. In general, full epistatic landscapes are intractable to compute for design spaces of practical sizes, but estimation of a subset of the terms is an active area of research (e.g.otwinowski13_truncated-epistatic-regression, park24_rfa) due to beliefs that natural proteins tend to exhibit sparse and mostly lower-order epistasis.
In comparison to the function decompositions used by DADO, a full epistatic landscape specifies both the graph topology and the component functions, everywhere. Completely specifying an order-$d$ interaction if each variable has $D$ states requires choosing $D^{d}$ values. The high-dimensional synthetic functions used in our experiments are fully specified everywhere because they’re defined with primarily lower-order epistasis, which can be specified with far fewer values. In contrast, our protein property predictors can have high-order component functions represented by neural networks fit from limited data, meaning they’re underspecified beyond the training data. In some sense, fitting neural network component functions allows us to specify higher-order epistasis in a locale without paying the price for defining it over the full design space. Real-world design procedures typically only have knowledge of $f$ in a small locale, which is why locally-valid decompositions suffice.
A.2 Decomposition-unaware EDA derivation
Given an objective, $f(x)$, to maximize, a decomposition-unaware EDA (or “naive” EDA) transforms the original problem into a distributional optimization problem as follows [brookes20_eda_em]:
$\max_{x}f(x)=\max_{\theta}\mathbb{E}_{p_{\theta}(x)}[f(x)].$ (16)
For the equivalence to hold, the search distribution, $p_{\theta}(x)$, must be capable of representing a point mass on $x^{*}$. Intuitively, one can think of an EDA as having a search distribution, $p_{\theta}(x)$, that acts as a spotlight on the design space, which is iteratively moved toward areas of the space with high $f(x)$ in expectation. Modern EDAs parameterize the search distribution with neural network generative models, such as a Variational Autoencoder (kingma13_vae; e.g.,  brookes19_cbas).
Procedurally, after having initialized the search distribution, $p_{\theta^{0}}(x)$, this naive EDA then iterates through these three steps either $N$ times or until some convergence criteria is met [larranaga01_eda, brookes20_eda_em]:
  1. 1.
    Draw $K$ samples from the current search distribution, $\{x^{k}\}_{k=1}^{K}\sim p_{\theta^{n}}(x)$
  2. 2.
    Score each sample with the objective function to get its weight, $w^{k}=f(x^{k})$.
  3. 3.
    Update the search distribution parameters with weighted maximum likelihood estimation (MLE), using the weighted samples, $\theta^{n+1}=\arg\max_{\theta}\;\mathbb{E}_{\{x^{k}\}}[w^{k}\log p_{\theta}(x^{k})$]. In older EDAs, this step was often instead a truncated maximum likelihood estimation, in which $p_{\theta}(x)$ modeled only samples with the largest weights.
Often, a predefined monotonic “shaping” function, $W(\cdot)$, is additionally applied to the weights to control optimization dynamics. Its monotonicity guarantees that $\arg\max_{x}f(x)=\arg\max_{x}W(f(x))$. We’ve written the EDA above and its derivation below without $W$, but $f(x)$ can be equivalently replaced with $W(f(x))$ everywhere.
For clarity, we sketch out one way of deriving the EDA update (step 3) from Eq. 16. At each EDA iteration, we want to improve the current search distribution with an update, $p_{\theta^{n+1}}(x)\leftarrow p^{*}_{n}(x)\propto f(x)\cdot p_{\theta^{n}}(x)$. To do this, we maximize
$-D_{\mathrm{KL}}\!\left(p^{*}_{n}(x)\,\|\,p_{\theta}(x)\right)=\frac{1}{Z}\mathbb{E}_{p_{\theta^{n}}(x)}[f(x)\log p_{\theta}(x)]+H(p^{*}_{n}(x)),$ (17)
where the entropy term can be dropped because it has no dependence on $\theta$ and division by $Z$ can be dropped without changing the objective’s maximizer. The sample approximation of the remaining terms is exactly step 3. Assuming an exact update (i.e., infinite samples such that the expectation is evaluated exactly, $p_{\theta}(x)$ has sufficient capacity to represent $p^{*}_{n}(x)$, and the KL divergence reaches 0), the resulting search distribution at iteration $n$ is $p^{*}_{n}(x)\propto f(x)^{n}p_{\theta^{0}}(x)$. As $n\rightarrow\infty$, this distribution concentrates all of its mass at the global maxima of $f$ and as a result, $f(x^{*})=\mathbb{E}_{p_{\theta}(x)}[f(x)]$. The derivation in brookes18_dbas is perhaps closest to this one, though their search distribution improvement operator is motivated by Bayes’ rule. It can be generalized to the derivation we give by simply writing $p^{*}_{n}(x)\propto C(x)\cdot p_{\theta^{n}}(x)$ for some $C$. Bayes’ rule yields a special case of this improvement operator, namely one where the multiplier is $C(x)=\frac{p(S\mid x)}{\sum_{x}p_{\theta^{n}}(x)p(S\mid x)}$. We choose not to interpret $f(x)$ as a normalized distribution in our derivation, though one can definitely do so.
In practice, the weighted MLE problem in step 3 is typically solved with gradient descent; we will choose to use only a fixed number of gradient steps, which can be theoretically justified through the equivalence of EDAs to Expectation-Maximization [brookes20_eda_em].
A.3 DADO derivation
Herein, we give an augmented version of the derivation presented in Sec. 2.1. We build on the derivation of the naive EDA given in Sec. A.2. The goal of DADO is to obtain a generative model, $p_{\theta}(x)$, that maximizes the EDA objective, $\arg\max_{\theta}\mathbb{E}_{p_{\theta}(x)}[f(x)]$, while leveraging decomposability in $f(x)$ for optimization efficiency. DADO requires as input a decomposed version of $f(x)$, which can be described by an undirected junction tree, $\mathcal{T}:=(\mathcal{N},\mathcal{E})$, with nodes, $\mathcal{N}$, and edges, $\mathcal{E}$. As noted in the introduction and shown in our experiments on proteins, identifying useful decomposability is feasible in practice. Given the junction tree decomposition, we write $f(x)=\sum_{i\in\mathcal{N}}f_{i}(\tilde{x}_{i})+\sum_{(i,j)\in\mathcal{E}}f_{i,j}(\tilde{x}_{i},\tilde{x}_{j})$, where we refer to $f_{i}(\tilde{x}_{i})$ as “node component functions” and $f_{i,j}(\tilde{x}_{i},\tilde{x}_{j})$ as “edge component functions” (Fig. 1a). This functional structure is intimately related to notions of the “epistatic landscape” of a protein’s property function and spectral analysis (see Sec. A.1.1). When the component functions are not known a priori, they can be parameterized and fit to labeled data.
We will begin our exposition by recalling how to do decomposition-aware exact (non-distributional) optimization, that is, to solve $\arg\max_{x}f(x)$. This problem is efficiently solved with a classical message-passing algorithm, which coordinates local optimizations across parts of the junction tree to obtain a single global optimum. Its efficiency comes from breaking optimization over all variables jointly into separate optimizations for each (smaller) meta-variable. Having loaded the reader with this intuition, we then adapt these ideas to distributional optimization, yielding DADO.
A.3.1 Classical message-passing for non-distributional optimization
Although classical message-passing has been largely used for probabilistic inference on probabilistic graphical models [pearl88_bp_textbook, shah_14_mit_inference-algorithms_message-passing], it can also be used for exact optimization of a function, $x^{*}=\arg\max\nolimits_{x}f(x)$ [vlassis04_maxplus]. In particular, message-passing can be used to find a global maximum of a function defined on an undirected junction tree, $\mathcal{T}$, by making use of its topology for optimal time-complexity. In particular, one takes the junction tree, which is undirected, and roots it to obtain a directed tree, which induces a hierarchy among the meta-variables from root to leaves. Each node is responsible for accumulating information from all nodes in its sub-tree and then passing this information on to its parent. Consequently, the root node receives information from the entire tree, which is sufficient to set its variables in a globally optimal manner. Then, starting with the root, each parent communicates its variables’ optimal settings to its children, which can in turn set their variables optimally, and so forth.
To obtain the rooted tree, $\mathcal{T^{\prime}}:=(\mathcal{N^{\prime}},\mathcal{E^{\prime}})$, from the junction tree, one keeps the same nodes, $\mathcal{N^{\prime}}=\mathcal{N}$, chooses a root node, $r$, and directs all edges in $\mathcal{E}$ outward from $r$, yielding directed edges, $\mathcal{E^{\prime}}$. Although rooting at any node will suffice, we choose $r$ such that $\mathcal{T^{\prime}}$ has the shortest height possible. Message-passing finds a global optimum in two passes through the tree: one round of passing messages up from leaves to root, and then one round passing messages back down to the leaves.
Given $\mathcal{T^{\prime}}$, classical message-passing first uses dynamic programming to accumulate information from the leaves up to the root. Similarly to any dynamic programming procedure, we accumulate solutions to increasingly larger intermediate sub-problems. In this case, each sub-problem is to find the value of $f(x)$ evaluated on only a subset of meta-variables, rather than on the full set of variables in $x$. Each sub-problem is tractable owing to the decomposition of the objective function into component functions for each node and edge, and by respecting the partial order of sub-problems induced by $\mathcal{T^{\prime}}$. Specifically, one computes a value function, $V^{\text{max}}_{i}(\tilde{x}_{\text{p}(i)})$, for each node $i\in\mathcal{N^{\prime}}\setminus\{r\}$, which tells us for each setting of its parent, $\tilde{x}_{\text{p}(i)}$, the value of the intermediate objective function defined by the edge component function, $f_{\text{p}(i),i}(\tilde{x}_{\text{p}(i)},\tilde{x}_{i})$, plus all component functions over the sub-tree rooted at $i$, given that all nodes maximize their respective intermediate objectives. Computing value functions comprises the first pass through the tree, from leaves to root:
$V^{\text{max}}_{i}(\tilde{x}_{\text{p}(i)}):=\max\nolimits_{\tilde{x}_{i}}\Big(f_{i}(\tilde{x}_{i})+f_{\text{p}(i),i}(\tilde{x}_{\text{p}(i)},\tilde{x}_{i})+\sum\nolimits_{c\in\text{children}(i)}V^{\text{max}}_{c}(\tilde{x}_{i})\Big).$
Notably, $V^{\text{max}}_{i}(\tilde{x}_{\text{p}(i)})$ provides sufficient information about all nodes in the sub-tree rooted at $i$ to optimally choose the value of $\tilde{x}_{\text{p}(i)}$ with respect to its children. Thus it follows that once all value functions have been computed, the root’s assignment can be set in a globally optimal manner from its children’s value functions,
$\tilde{x}_{r}^{*}:=\arg\max\nolimits_{\tilde{x}_{r}}\Big(f_{r}(\tilde{x}_{r})+\sum\nolimits_{c\in\text{children}(r)}V_{c}(\tilde{x}_{r})\Big).$
Having chosen the root assignment, we then pass it down the tree as $\tilde{x}_{\text{p}(i)}=\tilde{x}_{\text{p}(i)}^{*}$ to its children, which successively pass their chosen assignments to their children, all the way to the leaves,
$\tilde{x}_{i}^{*}:=\arg\max\nolimits_{\tilde{x}_{i}}\Big(f_{i}(\tilde{x}_{i})+f_{\text{p}(i),i}(\tilde{x}_{\text{p}(i)}^{*},\tilde{x}_{i})+\sum\nolimits_{c\in\text{children}(i)}V_{c}(\tilde{x}_{i})\Big),$
resulting in a global maximizer $x^{*}$ of $f(x)$. This dynamic programming “traceback” of optimal assignments back down the tree constitutes our second and final pass of messages.
Alternative notation.
For convenience of our generalization to distributional optimization, we re-write the parent value functions $V^{\text{max}}_{i}(\tilde{x}_{\text{p}(i)})$ in terms of child-parent, $Q^{\text{max}}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})$, and single-node, $Q^{\text{max}}_{i}(\tilde{x}_{i})$ value functions:
$V^{\text{max}}_{i}(\tilde{x}_{\text{p}(i)})$ $:=\max\nolimits_{\tilde{x}_{i}}\;Q^{\text{max}}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)}),\text{\;\;where}$ (18)
$Q^{\text{max}}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})$ $:=f_{\text{p}(i),i}(\tilde{x}_{\text{p}(i)},\tilde{x}_{i})+Q^{\text{max}}_{i}(\tilde{x}_{i})\text{\;\;and}$ (19)
$Q^{\text{max}}_{i}(\tilde{x}_{i})$ $:=f_{i}(\tilde{x}_{i})+\sum\nolimits_{c\in\text{children}(i)}V^{\text{max}}_{c}(\tilde{x}_{i}).$ (20)
In contrast to the original parent value functions, $Q^{\text{max}}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})$ represents the effect of the choice of both $\tilde{x}_{\text{p}(i)}$ and $\tilde{x}_{i}$ on their edge component function plus all component functions over the sub-tree rooted at $i$, assuming all descendants of $i$ maximize their corresponding value functions. Intuitively, $Q^{\text{max}}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})$ is the value function on edge $(\text{p}(i),i)$ prior to $\tilde{x}_{i}$ being maximized out, which will become useful if we want to, say, sample $\tilde{x}_{i}$ according to some distribution instead. $Q^{\text{max}}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})$ is composed of two terms, one of which depends on its parent, and one of which doesn’t, $Q^{\text{max}}_{i}(\tilde{x}_{i})$. Written using the $Q$-functions just defined, the equivalent traceback equations for selecting a globally optimal assignment, $x^{*}$, are simply $\tilde{x}_{r}^{*}:=\arg\max_{\tilde{x}_{r}}Q^{\text{max}}_{r}(\tilde{x}_{r})$ and $\tilde{x}_{i}^{*}:=\arg\max_{\tilde{x}_{i}}Q^{\text{max}}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)}^{*})$. In other words,
$x^{*}=\arg\max\nolimits_{x}f(x)$ $=\{\arg\max\nolimits_{\tilde{x}_{r}}Q^{\text{max}}_{r}(\tilde{x}_{r})\}\;+\;\{\arg\max\nolimits_{\tilde{x}_{i}}Q^{\text{max}}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)}^{*})\}_{(\text{p}(i),i)\in\mathcal{E^{\prime}}}$ (21)
$=\arg\max_{x}\Big(Q^{\text{max}}_{r}(\tilde{x}_{r})+\sum\nolimits_{(\text{p}(i),i)\in\mathcal{E^{\prime}}}Q^{\text{max}}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})\Big).$ (22)
A.3.2 From classical message passing to distributional optimization
In the same way that an EDA transforms $\arg\max_{x}f(x)$ into a distributional optimization problem (Sec. A.2), we can rewrite the equivalent message-passing objective in Eq. 22 as DO. Because the original optimization problems over $x$ are equivalent, their DO formulations are equivalent too,
$\arg\max\nolimits_{\theta}\mathbb{E}_{p_{\theta}(x)}[f(x)]$ $=\arg\max_{\theta}\mathbb{E}_{p_{\theta}(x)}\Big[Q^{\text{max}}_{r}(\tilde{x}_{r})+\sum\nolimits_{(\text{p}(i),i)\in\mathcal{E^{\prime}}}Q^{\text{max}}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})\Big]$ (23)
$=\arg\max_{\theta}\Big(\mathbb{E}_{p_{\theta}(x)}[Q^{\text{max}}_{r}(\tilde{x}_{r})]+\sum\nolimits_{(\text{p}(i),i)\in\mathcal{E^{\prime}}}\mathbb{E}_{p_{\theta}(x)}[Q^{\text{max}}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})]\Big),$ (24)
where we’ve used linearity of expectations. However, a generic joint search distribution cannot take advantage of the linear additivity in value functions over the junction tree topology. That is, while the classical traceback equations perform maximization over each meta-variable separately, Eq. 24 uses a single, unfactorized search distribution over all variables, $p_{\theta}(x)$, to optimize each $Q$-function. We address this next, by factorizing the search distribution.
Factorized search distribution.
Classical message-passing (Eq. 21) independently maximizes each $Q$-function conditional on the choice of $\tilde{x}_{\text{p}(i)}$, instead of explicitly maximizing all variables in $x$ jointly (Eq. 22). This is possible because each $Q_{i}$ captures all relevant global information needed to choose $\tilde{x}_{i}$. It stands to reason then that DO can do something similar. Specifically, instead of training a single joint search distribution, we train smaller search distributions over each $\tilde{x}_{i}$, conditional on $\tilde{x}_{\text{p}(i)}$, to separately maximize each $Q^{\text{max}}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})$. That is, we factor the search distribution according to $\mathcal{T^{\prime}}$, resulting in a DAG search distribution, $p_{\theta}(x)\coloneqq p_{\theta}(\tilde{x}_{r})\prod_{(\text{p}(i),i)\in\mathcal{E^{\prime}}}p_{\theta}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})$, for root node $r$, non-root nodes $i$, and parents $\text{p}(i)$, connected by directed edges $\mathcal{E^{\prime}}$, and with parameters, $\theta$. This factorized search distribution can be plugged into Eq. 24 for an equivalent optimization problem. We refer to each element of this product as one of the factors of the search distribution. In our implementation, each factor has completely separate parameters, though we write a shared $\theta$ for conciseness. Notice that each factor of the search distribution interacts with each other factor through the directed edges, $\mathcal{E^{\prime}}$. That is, the distribution of $\tilde{x}_{i}$ depends on its parent’s factor, and through it, all of its parent’s ancestors: $p_{\theta}(\tilde{x}_{i})=p_{\theta}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})p_{\theta}(\tilde{x}_{\text{p}(i)})$. In turn, each of node $i$’s children’s factors depends on $p_{\theta}(\tilde{x}_{i})$. Due to this coupling, we cannot optimize each factor fully independently. But we can still update each factor separately from the others in a globally-consistent manner via message-passing. In particular, each factor will only be responsible for directly optimizing its own meta-variable, but will need to coordinate with its neighboring factors by getting information from them about how they are optimizing their meta-variables in a manner analogous to classical message-passing. Our current messages, $Q_{i}^{\text{max}}$, convey the value of each intermediate objective when all meta-variables are chosen via maximization. For DO, we’ll require messages that communicate values when meta-variables are chosen according to the factorized search distribution, $p_{\theta}(x)$.
Distributional value functions.
While one certainly could choose to optimize classical value functions using an EDA search distribution (Eq. 24), it doesn’t make sense for two main reasons. As we just mentioned, DO aims to train $p_{\theta}(x)$ such that it maximizes $f(x)$, or equivalently, the sum of $Q$-functions, in expectation. Therefore, the DO objective should consider intermediate objective values for meta-variables chosen according to the current search distribution, not those chosen by explicit maximization, as in $V^{\text{max}}_{i}(\tilde{x}_{\text{p}(i)})$. Additionally, computing classical value functions requires enumerating all assignments of each $\tilde{x}_{i}$ for the maximum used in $V^{\text{max}}_{i}(\tilde{x}_{\text{p}(i)})$. When $\tilde{x}_{i}$ contains more than a few design variables, this max operation quickly becomes intractable. One reason for doing distributional optimization is to avoid enumerating massive design spaces. To address both issues, we define corresponding distributional value functions that fulfill both desiderata:
$V^{\theta}_{i}(\tilde{x}_{\text{p}(i)})$ $:=\mathbb{E}_{p_{\theta}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})}[Q^{\theta}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})],\text{\;\;where}$ (25)
$Q^{\theta}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})$ $:=f_{\text{p}(i),i}(\tilde{x}_{\text{p}(i)},\tilde{x}_{i})+Q^{\theta}_{i}(\tilde{x}_{i})\text{\;\;and}$ (26)
$Q^{\theta}_{i}(\tilde{x}_{i})$ $:=f_{i}(\tilde{x}_{i})+\sum\nolimits_{c\in\text{children}(i)}V^{\theta}_{c}(\tilde{x}_{i}).$ (27)
Compared to Eq. 18, the distributional $V$-functions compute the value in expectation under $p_{\theta}$ instead of a $\max$. Because they use an expectation instead of a $\max$ operation, these value functions can be approximated tractably and without bias by drawing Monte-Carlo samples from the search distribution. The connection to classical value functions can be made clearer by defining a particular distribution which recovers them, the one placing all of its mass on the $\arg\max$, and its corresponding distributional value functions:
$\quad\quad\quad\quad p_{\text{max}}(\tilde{x}_{i}=A\mid\tilde{x}_{\text{p}(i)})=\begin{cases}\;\;1&\quad\text{if }A=\arg\max_{\tilde{x}_{i}}Q^{\text{max}}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})\\ \;\;0&\quad\text{otherwise}\end{cases},$ (28)
$V^{\text{max}}_{i}(\tilde{x}_{\text{p}(i)})=\mathbb{E}_{p_{\text{max}}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})}[Q^{\text{max}}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})]=\max_{\tilde{x}_{i}}f_{i}(\tilde{x}_{i})+f_{\text{p}(i),i}(\tilde{x}_{\text{p}(i)},\tilde{x}_{i})+\sum_{c\in\text{children}(i)}V^{\text{max}}_{c}(\tilde{x}_{i}),$ (29)
$Q^{\text{max}}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})=f_{\text{p}(i),i}(\tilde{x}_{\text{p}(i)},\tilde{x}_{i})+Q^{\text{max}}_{i}(\tilde{x}_{i})\text{, \;\;\;for }Q^{\text{max}}_{i}(\tilde{x}_{i})=f_{i}(\tilde{x}_{i})+\sum_{c\in\text{children}(i)}V^{\text{max}}_{c}(\tilde{x}_{i})$ (30)
Moreover, each distributional value function lower bounds each corresponding classical value function because the expectation of a function cannot exceed its maximum. This relation is evident for the base case of a leaf node,
$V^{\text{max}}_{i}(\tilde{x}_{\text{p}(i)})=\max_{\tilde{x}_{i}}f_{i}(\tilde{x}_{i})+f_{\text{p}(i),i}(\tilde{x}_{\text{p}(i)},\tilde{x}_{i})\geq\mathbb{E}_{p_{\theta}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})}[f_{i}(\tilde{x}_{i})+f_{\text{p}(i),i}(\tilde{x}_{\text{p}(i)},\tilde{x}_{i})]=V^{\theta}_{i}(\tilde{x}_{\text{p}(i)}),$ (31)
and by inductive argument, is also true for all other nodes’ $V$-functions, and similarly, the $Q$-functions, which only differ in which $V$-functions they sum over. This implies that the sum of classical value functions in the objective in Eq. 24 is bounded below by the sum of distributional value functions. This bound still holds when we take the expectation of each sum under $p_{\theta}(x)$,
$\mathbb{E}_{p_{\theta}(x)}[Q^{\text{max}}_{r}(\tilde{x}_{r})]+\sum_{(\text{p}(i),i)\in\mathcal{E^{\prime}}}\mathbb{E}_{p_{\theta}(x)}[Q^{\text{max}}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})]\geq\mathbb{E}_{p_{\theta}(x)}[Q^{\theta}_{r}(\tilde{x}_{r})]+\sum_{(\text{p}(i),i)\in\mathcal{E^{\prime}}}\mathbb{E}_{p_{\theta}(x)}[Q^{\theta}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})].$ (32)
We’ll optimize the lower bound with its distributional value functions instead of the sum of classical value functions.
Distributional optimization with value functions.
All that remains is to derive an update rule that treats each search distribution factor separately. Since each expectand in Eq. 32 doesn’t depend on descendant meta-variables, we can replace $p_{\theta}(x)$ with each meta-variable’s marginal distribution:
$\arg\max_{\theta}\mathbb{E}_{p_{\theta}(\tilde{x}_{r})}[Q^{\theta}_{r}(\tilde{x}_{r})]+\sum\nolimits_{(\text{p}(i),i)\in\mathcal{E^{\prime}}}\mathbb{E}_{p_{\theta}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})p_{\theta}(\tilde{x}_{\text{p}(i)})}[Q^{\theta}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})].$ (33)
We then follow the EDA derivation (Sec. A.2) to arrive at a sample-approximated update rule using the previous iteration’s distribution, $p_{\theta^{n}}(x)$ in which each term is optimized by only a single factor. We omit the derivation for the root factor because it’ss exactly the same as in Sec. A.2, using $Q_{r}^{\theta^{n}}(\tilde{x}_{r})$ as the weight instead of $f(x)$. To update a conditional factor from Eq. 33 in an EDA loop, we define $p^{*}_{n}(\tilde{x}_{c}\mid\tilde{x}_{\text{p}(i)})\propto Q_{c}^{\theta^{n}}(\tilde{x}_{c},\tilde{x}_{\text{p}(i)})\cdot p_{\theta^{n}}(\tilde{x}_{c}\mid\tilde{x}_{\text{p}(i)})$, and minimize its divergence from $p^{\theta}(\tilde{x}_{c}\mid\tilde{x}_{\text{p}(i)})$ using the law of iterated expectations:
$\quad\quad\quad\quad\quad\quad-\mathbb{E}_{p_{\theta^{n}}(\tilde{x}_{\text{p}(i)})}\left[D_{\mathrm{KL}}\!\left(p^{*}_{n}(\tilde{x}_{c}\mid\tilde{x}_{\text{p}(i)})\,\|\,p_{\theta}(\tilde{x}_{c}\mid\tilde{x}_{\text{p}(i)})\right)\right]$ (34)
$=\;$ $\frac{1}{Z}\mathbb{E}_{p_{\theta^{n}}(\tilde{x}_{\text{p}(i)})}\left[\mathbb{E}_{p_{\theta^{n}}(\tilde{x}_{c}\mid\tilde{x}_{\text{p}(i)})}[Q_{c}^{\theta^{n}}(\tilde{x}_{c},\tilde{x}_{\text{p}(i)})\log p_{\theta}(\tilde{x}_{c}\mid\tilde{x}_{\text{p}(i)})]+H(p^{*}_{n}(\tilde{x}_{c}\mid\tilde{x}_{\text{p}(i)}))\right].$ (35)
Again, when taking the $\arg\max$ with respect to $\theta$, we can equivalently remove the entropy term which bears no dependence on $\theta$ and division by constant $Z$. Notice that the dependence of the conditional terms on $p_{\theta}(\tilde{x}_{\text{p}(i)})$ in Eq. 33 is approximated by sampling, which is why only individual factors of the search distribution, $\log p_{\theta}(\tilde{x}_{c}\mid\tilde{x}_{\text{p}(i)})$, appear in each summand, as opposed to $\log p_{\theta}(\tilde{x}_{c}\mid\tilde{x}_{\text{p}(i)})p_{\theta}(\tilde{x}_{\text{p}(i)})$. Updating $p_{\theta}(\tilde{x}_{\text{p}(i)})$ to maximize $Q_{c}^{\theta}(\tilde{x}_{c},\tilde{x}_{\text{p}(i)})$ would be redundant and against the spirit of message-passing since parent $Q$-functions already include that info from children. The resulting update rule over all search distribution factors (sharing a single set of samples; see Fig. 1b) can be written as a sum of weighted likelihoods for each factor,
$\theta^{n+1}=\arg\max_{\theta}\sum\nolimits_{x\sim p_{\theta^{n}}(x)}\Big(Q_{r}^{\theta^{n}}(\tilde{x}_{r})\log p_{\theta}(\tilde{x}_{r})+\sum\nolimits_{(\text{p}(i),i)\in\mathcal{E^{\prime}}}Q_{i}^{\theta^{n}}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})\log p_{\theta}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})\Big),$ (36)
which can be written equivalently as separate updates because the factors don’t share parameters, as desired:
$\theta_{r}^{n+1}$ $=\arg\max_{\theta_{r}}\sum\nolimits_{x\sim p_{\theta^{n}}(x)}Q_{r}^{\theta^{n}}(\tilde{x}_{r})\log p_{\theta_{r}}(\tilde{x}_{r})\text{, \;\; and}$ (37)
$\theta_{i}^{n+1}$ $=\arg\max_{\theta_{i}}\sum\nolimits_{x\sim p_{\theta^{n}}(x)}Q_{i}^{\theta^{n}}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})\log p_{\theta_{i}}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)}),\;\;\forall\;i\in\mathcal{N}\setminus\{r\}.$ (38)
Each factor is weighted by its corresponding value function, enabling it to coordinate with all its descendant factors despite their being updated separately. The whole DO is tied together at the top by the root factor. Our resulting algorithm (Alg. 1) fits into the three EDA steps: (1) designs are sampled from $p_{\theta^{n}}(x)$, (2) weights, here each $Q$-function instead of $f(x)$, are computed, and (3), each search distribution factor receives its own independent weighted maximum likelihood update based on these weights (steps 1 and 2 illustrated in Fig. 1b). These are repeated until convergence, or for some fixed number of iterations. The factor updates are coupled only through the $Q$-functions, the messages across edges. $\{Q^{\theta^{n}}_{i}\}_{i\in\mathcal{N}}$ are only valid while $\theta$ is close to $\theta^{n}$, meaning one must balance how many gradient steps are taken before drawing new samples. If too many gradient steps are taken without resampling, a factor may be changing its parameters to collaborate with another factor which has already changed its behavior. It’s common for EDAs to include an additional hyperparameter $W(\cdot)$, a monotonic shaping function applied to the weights, to alter optimization dynamics [brookes20_eda_em], so we include it in Alg.1( 1 on lines 16 and 18. Choosing $W$ to be the identity recovers our derivation above. Notice that in place of the classical message-passing traceback equations, DADO simply performs sequential conditional sampling from its search distribution. Interestingly, a very similar algorithm can be derived without using a lower bound on the value functions (Eq. 32) if one adds an entropy-maximizing term to the initial DO objective (Sec. A.4).
A.4 Alternate DADO Derivation: Maximum Entropy Objective
If, by comparison to Sec. A.3, we instead seek to solve a closely related objective which additionally maximizes the search distribution’s entropy, we can arrive at a similar method by leveraging the probabilistic graphical modeling (PGM) framework in levine18_control-as-inference. This view will provide another intuition for the relationship between classical value functions and distributional value functions approximated with Monte-Carlo samples. It also highlights the core components of our conceptual framework, which are shared even by derivations starting from different objectives. In other words, one need not use our specific method as outlined in Alg. 1 to reap the benefits of decomposition-aware distributional optimization, and variations may have desirable properties and/or simply perform better in certain scientific design settings. This derivation will assume that you’ve already read key parts of Sec. 2A.3 and are familiar with the function decomposition, search distribution factorization, and value functions.
Maximum entropy decomposition-aware distributional optimization.
We begin from the maximum entropy problem [jaynes57_mep, ziebart08_maxent, zhu24_aav_max-entropy],
$\arg\max_{\theta}\mathbb{E}_{p_{\theta}(x)}[f(x)]+\beta H(p_{\theta}),\;\;\beta\geq 0.$ (39)
This objective is arguably more consistent with the desired end result for many scientific design problems. Often, a distribution over good solutions is preferred to a single solution (i.e., a distribution with no entropy) [^1] In our experiments (no entropy bonus), the search distribution avoids collapse because we don’t run until convergence (fixed number of iterations), and because we sweep the learning rate (if it’s very high, the distribution may quickly collapse to a point, limiting further improvement). Similar strategies were used by brookes19_cbas instead of an explicit entropy bonus. EDAs were originally used for discrete optimization problems where a single solution was desired. Using distributional optimization to solve more nuanced problems calls for some modifications.. When optimizing a predictive model or more broadly, dealing with uncertainty or inaccuracy in the objective function, having a distribution over good solutions can be particularly important. The maximum entropy objective has a closed-form solution, $p^{*}(x)\propto\exp f(x)/\beta$, motivating the solution of an equivalent variational objective, $\arg\min_{\theta}D_{\mathrm{KL}}\!\left(p_{\theta}\,\|\,p^{*}\right)$, from which we will obtain a decomposed update rule. We first plug in decomposed versions of $f$ and $p_{\theta}(x)$, as described in Sec.Sec. 2.1A.3, to get
$\arg\min_{\theta}D_{\mathrm{KL}}\!\left(p_{\theta}\,\|\,p^{*}\right)=\arg\min_{\theta}D_{\mathrm{KL}}\!\left(p_{\theta}(\tilde{x}_{r})\prod_{i\in\mathcal{N}\setminus\{r\}}p_{\theta}(x_{i}\mid\tilde{x}_{\text{p}(i)})\,\|\,\exp\frac{\sum_{i\in\mathcal{N}}f_{i,p}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})+f_{i}(\tilde{x}_{i})}{\beta}\right).$ (40)
Notice that we’ve dropped the normalizing constant for $p^{*}$ because it has no dependence on $\theta$. In what follows, for compactness and clarity, we will abuse notation slightly by including the root factor, $p_{\theta}(\tilde{x}_{r})$, with the other nodes, writing it as conditional on a parent even though it has no parent and is an unconditional distribution.
$-D_{\mathrm{KL}}\!\left(\prod_{i\in\mathcal{N}}p_{\theta}(x_{i}\mid\tilde{x}_{\text{p}(i)})\,\|\,\exp\frac{\sum_{i\in\mathcal{N}}f_{i,p}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})+f_{i}(\tilde{x}_{i})}{\beta}\right)$ (41)
$=\mathbb{E}_{p_{\theta}(x)}\left[\frac{\sum_{i\in\mathcal{N}}f_{i,p}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})+f_{i}(\tilde{x}_{\text{p}(i)})}{\beta}-\log\prod_{i\in\mathcal{N}}p_{\theta}(x_{i}\mid\tilde{x}_{\text{p}(i)})\right]$ (42)
$=\mathbb{E}_{p_{\theta}(x)}\left[\sum_{i\in\mathcal{N}}\frac{f_{i,p}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})+f_{i}(\tilde{x}_{\text{p}(i)})}{\beta}-\log p_{\theta}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})\right]$ (43)
$=\sum_{i\in\mathcal{N}}\mathbb{E}_{p_{\theta}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})}\left[\frac{f_{i,p}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})+f_{i}(\tilde{x}_{\text{p}(i)})}{\beta}-\log p_{\theta}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})\right]$ (44)
The final line decomposes over nodes, but still has an explicit dependence on all ancestor node search distribution factors. One way to obtain a separate update for each search distribution factor is to derive a form of $p^{*}(x)$ factorized in accordance with $p_{\theta}(x)$’s factorization, so that we can simply minimize each factor’s divergence from the optimal one.
Optimal search distribution factors via message-passing.
To derive the optimal factors, $p^{*}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})$, we’ll use message-passing, for which we now introduce some notation. Note that we adopt the same convention as above of including the root with the rest of the value functions, meaning writing it with a parent, although it doesn’t have a parent and all terms involving its nonexistent parent are simply 0. We recursively define value functions (consistent with levine18_control-as-inference) as:
$\quad\quad\quad\quad\quad\quad\quad\quad\quad V_{i}(\tilde{x}_{\text{p}(i)})=\log\sum\nolimits_{\tilde{x}_{i}}\exp Q_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)}),$ (45)
$Q_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})=f_{\text{p}(i),i}(\tilde{x}_{\text{p}(i)},\tilde{x}_{i})+Q_{i}(\tilde{x}_{i})\text{, and }Q_{i}(\tilde{x}_{i})=f_{i}(\tilde{x}_{i})+\sum\nolimits_{c\in\text{children}(i)}V_{c}(\tilde{x}_{i}).$ (46)
Though they might seem a bit arbitrary, substituting the value functions into Eq. 44 will help illuminate $p^{*}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})$. Notice that the $Q$-functions have the same recursive definition as in DADO, though the $V$-functions are different. We first consider the base case of a leaf node, where $Q_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})=f_{\text{p}(i),i}(\tilde{x}_{\text{p}(i)},\tilde{x}_{i})+f_{i}(\tilde{x}_{i})$. We substitute, add $0=V_{i}(\tilde{x}_{\text{p}(i)})/\beta-V_{i}(\tilde{x}_{\text{p}(i)})/\beta$, and use the definition of KL divergence:
$\mathbb{E}_{p_{\theta}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})p_{\theta}(\tilde{x}_{\text{p}(i)})}\left[\log\exp\frac{Q_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})}{\beta}-\log p_{\theta}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})+\frac{V_{i}(\tilde{x}_{\text{p}(i)})}{\beta}-\log\exp\frac{V_{i}(\tilde{x}_{\text{p}(i)})}{\beta}\right]$ (47)
$=\;$ $\mathbb{E}_{p_{\theta}(\tilde{x}_{\text{p}(i)})}\left[-D_{\mathrm{KL}}\!\left(p_{\theta}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})\,\|\,\exp\frac{Q_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})-V_{i}(\tilde{x}_{\text{p}(i)})}{\beta}\right)+\frac{V_{i}(\tilde{x}_{\text{p}(i)})}{\beta}\right].$ (48)
Because $p_{\theta}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})$ only appears in the KL divergence, the overall expectation will be maximized, as far as $p_{\theta}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})$ is concerned, when the divergence is 0. Therefore, for a leaf node, $p^{*}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})=\exp((Q_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})-V_{i}(\tilde{x}_{\text{p}(i)}))/\beta)$, which is a proper distribution because $\exp V_{i}(\tilde{x}_{\text{p}(i)})$ is exactly the normalizing constant of $\exp Q_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})$. Notice that we still have a term, $\mathbb{E}_{p_{\theta}(\tilde{x}_{\text{p}(i)})}\left[{V_{i}(\tilde{x}_{\text{p}(i)})}/{\beta}\right]$, outside of the divergence (Eq. 48). Maximizing this term must therefore be the responsibility of $p_{\theta}(\tilde{x}_{\text{p}(i)})$, not $p_{\theta}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})$. In fact, these leftover terms are included in the parent’s $Q$-function via the sum over all child $V$-functions in $Q_{i}(\tilde{x}_{i})$ as we defined it above (Eq. 46). It follows that the recursive case has the same form as the leaves, which had $Q_{i}(\tilde{x}_{i})=f_{i}(\tilde{x}_{i})$ because they’ve no children (more detailed derivation in levine18_control-as-inference). That is, for all non-root nodes $i$, $p^{*}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})=\exp((Q_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})-V_{i}(\tilde{x}_{\text{p}(i)}))/\beta)$; for the root node, $p^{*}(\tilde{x}_{r})=\exp((Q_{r}(\tilde{x}_{r})-V_{r})/\beta)$, where $V_{r}=\log\sum_{\tilde{x}_{r}}\exp Q_{r}(\tilde{x}_{r})$ is a normalizing constant with no dependence on $\theta$. We now have an optimal form for each factor and can proceed to minimize our search factors’ divergences from them completely separately.
Distributional optimization by matching optimal factors.
Given a factorization of the optimal search distribution, $p^{*}(x)$, we can train $p_{\theta}(x)$’s factors to match their corresponding optimal factor distributions, resulting in an algorithm very similar to DADO. The objective for node $i$’s factor minimizes its divergence to the optimal factor $p^{*}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})$,
$\arg\min_{\theta_{i}}\mathbb{E}_{p_{\theta_{\text{p}(i)}}(\tilde{x}_{\text{p}(i)})}\left[D_{\mathrm{KL}}\!\left(p_{\theta_{i}}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})\,\|\,\exp\frac{Q_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})-V_{i}(\tilde{x}_{\text{p}(i)})}{\beta}\right)\right]$ (49)
$=\;$ $\arg\max_{\theta_{i}}\mathbb{E}_{p_{\theta_{\text{p}(i)}}(\tilde{x}_{\text{p}(i)})}\left[\mathbb{E}_{p_{\theta_{i}}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})}\left[Q_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})\right]-V_{i}(\tilde{x}_{\text{p}(i)})+\beta H(p_{\theta_{i}}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)}))\right],$ (50)
[…]
$\theta^{n+1}_{i}=\arg\max_{\theta_{i}}\sum_{x\sim p_{\theta^{n}}(x)}\biggl(\Bigl(Q_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})-V_{i}(\tilde{x}_{\text{p}(i)})\Bigr)\log p_{\theta_{i}}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})-\beta\log p_{\theta_{i}}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})\biggr),$ (51)
and since the weight should always be non-negative, a monotonic shaping function, $W$, outputting non-negative weights, can be introduced too,
$\theta^{n+1}_{i}=\arg\max_{\theta_{i}}\sum_{x\sim p_{\theta^{n}}(x)}\biggl(W\Bigl(Q_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})-V_{i}(\tilde{x}_{\text{p}(i)})-\beta\Bigr)\log p_{\theta_{i}}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})\biggr).$ (52)
There are a few minor differences from DADO. First, there’s an entropy-maximizing term. Second, notice that the value function definitions are a little different from in DADO. Here we use $\log\sum\exp$ to define $V_{i}(\tilde{x}_{\text{p}(i)})$, whereas DADO uses an expectation over the search distribution instead. Third, the $Q$-function has $V_{i}(\tilde{x}_{\text{p}(i)})$ subtracted from it, whereas DADO’s derivation doesn’t. As mentioned above though, $V_{i}(\tilde{x}_{\text{p}(i)})$’s role in the objective is that of a “baseline”, and it can be changed out or dropped. In our implementation, we do subtract $V_{i}(\tilde{x}_{\text{p}(i)})$—the expectation-based one, not $\log\sum\exp$—as a mean baseline function to reduce weight variance and obtain a more stable update (Sec. 4). Overall though, the result is quite similar to DADO (Sec. A.3), supporting the generality of the decomposition-aware distributional optimization framework, and suggesting that there are a variety of related objectives and algorithms one might use under this umbrella. In particular, one might adapt methods from RL. For example, peng19_awr’s derivation could be adapted to get a similar weighted maximum likelihood objective with the difference in value functions (also called the advantage) exponentiated instead. TRPO, PPO, or CbAS [schulman15_trpo, schulman17_ppo, brookes19_cbas] might be adapted to regularize the search distribution toward some (identically factorized) prior, or the previous iteration’s search distribution for stability. We emphasize that the core idea of DADO is to leverage message-passing (i.e., value functions) on arbitrarily decomposed objectives to derive a distributional optimization procedure which updates search distribution factors separately, regardless of the specific objective and algorithm used.
From classical value functions to distributional value functions.
Under this definition of value functions using $\log\sum\exp$ (Eq. 45), we can view $V_{i}(\tilde{x}_{\text{p}(i)})$ as a soft maximum, which becomes a hard maximum when $Q_{i}$ are large, such that for $\tilde{x}_{i}^{*}$, $V_{i}(\tilde{x}_{\text{p}(i)})\approx Q_{i}(\tilde{x}_{i}^{*},\tilde{x}_{\text{p}(i)})$. The hard maximum case resembles classical message-passing, from which we began. When $Q_{i}$ are relatively small, information about multiple $\tilde{x}_{i}$ will pass through $\log\sum\exp$, thus propagating info for a distribution of descendant states in accordance with $p^{*}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})\propto\exp Q_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})/\beta$. Another way of seeing the transition from classical message-passing to distributional optimization is that for $\beta$ near 0, $p^{*}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})$ will concentrate all its mass on the $\arg\max$. But for a larger $\beta$, $p^{*}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})$ will concentrate on multiple designs proportional to their $Q$ values. As $\beta\rightarrow\infty$, $p^{*}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})$ becomes a uniform distribution.
A.5 Protein Dataset Details
We chose four protein property datasets based on three being readily available in a convenient form from Mave2022, and one from notin23_proteingym that has sequences many mutations away from the wild-type assayed. These proteins span various lengths and all have more than single mutations away from wild-type. These datasets are:
  • Amyloid-beta (Amyloid; $L=42,D=21$, where the extra state is a stop codon), with data from seuma21_amyloid_beta, which assayed 16,066 sequences for aggregation with a nucleation score, 97% of which are double mutants. The dataset was accessed via Mave2022.
  • Adeno-associated virus 2 capsid protein (AAV2 capsid or AAV; $L=28,D=20$), with data from bryant21_aav, which assayed 42,329 sequences for virus viability, including sequences with as many as 27 mutations from the wild-type. The full protein is of length 735. The dataset was accessed via notin23_proteingym.
  • Protein G B1 domain (GB1; $L=55,D=20$), with data from olson14_gb1, who scanned all possible single and double mutants and assayed their 536,963 binding affinities to IgGFC, of which 99.8% are double mutants. The dataset was accessed via Mave2022.
  • TAR DNA-binding protein 43 (TDP-43; $L=84,D=21$, where the extra state is a stop codon), with data  bolognesi19_tdp43, which assayed 57,996 sequences for cell toxicity, 98% of which are double mutants. The dataset was accessed via Mave2022.
We considered a few other datasets only shown in experiments in the appendix (Sec. A.6.2). These were also chosen from  notin23_proteingym on the basis of containing sequences many mutations away from the wild-type assayed, plus one arbitrarily chosen from the shortest-length (for expediency) domains in  tsuboyama23_mega. These datasets are:
  • ynzC, a small protein domain ($L=39,D=20$), with data from tsuboyama23_mega, assayed for its folding stability on 2,301 variants, 68% of which are double mutants. The dataset was accessed via notin23_proteingym.
  • Chlamydomonas reinhardtii’s light-oxygen-voltage domain (CreiLOV or Phot; $L=118,D=20$), with data from chen23_phot, which assayed 167,530 sequences for fluorescence, including sequences with as many as 15 mutations from the wild-type. The dataset was accessed via notin23_proteingym.
  • Yeast transcription factor Gcn4 ($L=44,D=20$), with data from staller18_gcn4, which assayed 2,639 sequences for activity, including sequences with as many as 44 mutations from the wild-type. The full protein is of length 281. The dataset was accessed via notin23_proteingym.
It’s worth noting that the Gcn4 activation domain assayed is intrinsically disordered, but folds into a helix when it binds its coactivator [staller18_gcn4]. We expect that our rudimentary procedure of obtaining a contact map from an AlphaFold3 [abramson24_af3]) 3D structure won’t work for proteins that exist in more than one stable configuration because the resulting decomposition will not capture interactions crucial to the assayed property. We find that the resulting predictive model has very poor holdout accuracy. Developing more accurate decompositions for such problems is an interesting line of future work.
A.6 Additional experimental results
Here, we include additional experimental results for completeness. For synthetic landscapes and for protein landscapes, we ablate the number of samples drawn at each EDA iteration compared to the main text figures. For protein landscapes, we also perform a case study on one protein dataset in order to study the effect of the distance threshold, $t$, on decomposability, predictive accuracy, and downstream optimization performance.
A.6.1 Synthetic functions
We first present an ablation on number of samples drawn each iteration, $K$, between DADO and a decomposition-unaware EDA on synthetic functions, keeping everything else the same as in Sec. A.6.1. For this ablation, we only consider synthetic landscapes with sequence lengths of $L=25$, $L=50$, $L=100$; we omit $L=200$ (Fig. 2c) for expediency. Optimization results for both $K=100$ (Fig. A1) and $K=1000$ (Fig. A2) support the conclusion that DADO is able to find designs with higher $f(x)$ earlier than the EDA, and by the end of 100 iterations, has a comparable or larger sample mean than the EDA. Both methods, as one would expect when performing distributional optimization, achieve faster and more effective optimization when the sampling budget is increased. DADO shows small performance improvements when given tenfold samples, suggesting that it is pretty sample-efficient and has already reached saturating performance by $K=100$, whereas the EDA finds samples with substantially higher $f(x)$ when given the same increased sample budget, demonstrating that it had not yet saturated at $K=100$.
[…]
Figure A1: Comparison of a standard EDA to DADO on synthetic problems drawing 100 samples each iteration. We created three random functions, $f(x)$, each with a randomly chosen junction tree decomposition with maximum node size of one, and randomly chosen parameters. Each experiment used alphabet size $D=20$ and sequence length, a, $L=25$ (also Fig. 2a), b, $L=50$ (also Fig. 2b), and c, $L=100$. Each of the two methods drew $K=100$ samples per iteration. For each iteration, we show the mean (solid line) and 95% confidence interval (shaded envelope) of the 100 samples evaluated on $f(x)$, averaged across results from 20 random seeds. P-values are from two-sided paired t-tests that AUC of the per-iteration mean is different between methods, using the 20 mean curves.
[…]
Figure A2: Comparison of a standard EDA to DADO on synthetic problems drawing 1000 samples each iteration. We created three random functions, $f(x)$, each with a randomly chosen junction tree decomposition with maximum node size of one, and randomly chosen parameters. Each experiment used alphabet size $D=20$ and sequence length, a, $L=25$, b, $L=50$, and c, $L=100$. Each of the two methods drew $K=1000$ samples per iteration. For each iteration, we show the mean (solid line) and 95% confidence interval (shaded envelope) of the 100 samples evaluated on $f(x)$, averaged across results from 20 random seeds. P-values are from two-sided paired t-tests that AUC of the per-iteration mean is different between methods, using the 20 mean curves.
A.6.2 Protein property predictors
We now present an ablation on seven protein landscapes for the number of samples at each iteration, $K$, keeping all other settings the same as in Sec. 4.2. These include the four proteins originally shown in the main text—Amyloid, AAV, GB1, and TDP-43 (Fig. 3)—as well as three proteins not shown elsewhere—ynzC, CreiLOV, and Gcn4—so that all can be viewed simultaneously and compared. The datasets upon which each of these protein property predictors is based are described in Sec. A.5.
We first present experimental results on the seven synthetic landscapes with $K=1000$ (Fig. A3). For this setting, optimization performance for both DADO and the EDA is saturated and comparable for two of the additional proteins, ynzC and CreiLOV (Fig. A3e,f). On Gcn4, for which the decomposed predictive model is based on the structure of an intrinsically disordered protein and so does not fit the assay very well (details in Sec. A.5), DADO is able to find high $f(x)$ designs more effectively than the EDA. All other predictive models acheived good holdout accuracy; we include Gcn4 primarily as an example of a protein lacking a good decomposition.
[…]
Figure A3: Comparison of a standard EDA to DADO on protein property predictive models with $\mathbf{K=1000}$. For each of seven proteins of varying length, a, Amyloid (also Fig. 3a), b, AAV (also Fig. 3b), c, GB1 (also Fig. 3c), d, TDP-43 (also Fig. 3d), e, ynzC, f, CreiLOV, and g, Gcn4, we fit a neural network property function, $f(x)$, adhering to a junction tree decomposition derived from the protein’s 3D structure, and then used standard EDA and DADO to optimize them. Each approach drew $K=1000$ samples per EDA iteration. For each iteration, we show the mean (solid line) and 95% confidence interval (shaded envelope) of the 1000 samples evaluated on $f(x)$, averaged across results from 20 random seeds. P-values are from two-sided paired t-tests that AUC of the per-iteration mean is different between methods, using the 20 mean curves.
For comparison, we show the same seven synthetic landscapes when DADO and the EDA can draw only $K=100$ samples each iteration (Fig. A4). DADO performs substantially better than the EDA on Amyloid and AAV still. On GB1, the EDA performs about as well with $K=100$ and $K=1000$. DADO is comparable to the EDA for $K=100$, but finds designs with much higher $f(x)$ when $K=1000$. For TDP-43 and ynzC, DADO finds designs with higher $f(x)$ than the EDA when $K=100$, whereas at $K=1000$, the two methods perform indistinguishably well. Both methods are comparable on CreiLOV for $K=100$ still, and DADO finds much better designs than EDA on Gcn4. Overall, as in Sec. A.6.1, we see that increasing $K$ helps, or at least does not hurt, both DADO and the EDA on all protein landscapes considered.
[…]
Figure A4: Comparison of a standard EDA to DADO on protein property predictive models with $\mathbf{K=100}$. For each of seven proteins of varying length, a, Amyloid, b, AAV, c, GB1, d, TDP-43, e, ynzC, f, CreiLOV, and g, Gcn4, we fit a neural network property function, $f(x)$, adhering to a junction tree decomposition derived from the protein’s 3D structure, and then used standard EDA and DADO to optimize them. Each approach drew $K=100$ samples per EDA iteration. For each iteration, we show the mean (solid line) and 95% confidence interval (shaded envelope) of the 100 samples evaluated on $f(x)$, averaged across results from 20 random seeds. P-values are from two-sided paired t-tests that AUC of the per-iteration mean is different between methods, using the 20 mean curves.
A.6.3 Investigation of protein property predictive model decomposability
We also sought to investigate how changing the threshold that determines the complexity of the junction tree would affect both predictive performance and optimization performance. the experiments, shown for experiments in which we draw 100 samples at each iteration. Specifically, we varied the distance threshold, $t$, for which pairs of residues were considered contacting to explore the tradeoff between accuracy of the model and decomposability. The lower the value of $t$, the stricter the decomposition (the smaller the cardinality of the largest meta-variable); thus lower $t$ should give DADO a larger advantage over the standard EDA, but may be overly restrictive and yield a worse predictive model. We decided to do a case study of GB1 for $K=100$ samples at each iteration because in this setting, using the default threshold of $t=4.5\text{\AA }$ (Fig. A4c) leads to comparable performance between DADO and the EDA. We wanted to see if decomposing the model further would result in a function that’s easier for DADO to optimize without sacrificing accuracy. We find that the largest $t$ provides the highest holdout predictive accuracy, which is expected because the resulting decomposed model isn’t restricted. However, decreasing this distance down to $t=2.75\text{\AA }$ allows most of the predictive signal to remain while substantially reducing the complexity of the junction tree (Fig. A5a), enabling DADO to consistently converge on designs with high $f(x)$ within only 10 iterations (Fig. A5b), whereas for $t=4.5\text{\AA }$ and $t=9\text{\AA }$, DADO’s mean does not clearly converge even after 100 iterations, and its distribution remains dispersed as evidenced by wide shaded envelopes (Fig. A5c,d). Additionally, DADO definitively outperforms the decomposition-unaware EDA when $t=2.75\text{\AA }$ (Fig. A5b). When the predictive model is less decomposed, DADO is not as distinguishable from the EDA (Fig. A5c,d).
[…]
Figure A5: Investigation of GB1 predictive model decomposability. a, We varied the distance threshold, $t$, for which pairs of residues in protein GB1’s 3D structure are considered neighbors in the junction tree, using several values in $t\in[0\text{\AA },15\text{\AA }]$, so as to explore the tradeoff between accuracy of the model and decomposability. The lower the value of $t$, the stricter the decomposition (the smaller the cardinality of the largest meta-variable); thus lower $t$ should give DADO a larger relative advantage over the standard EDA, but may yield a worse predictive model from imposing a more restrictive functional structure. The Pearson correlation on a 10% holdout set remains quite good down to and including $t=2.75\text{\AA }$, a point at which DADO provides a statistically significant win over a standard EDA as seen in panel b. Blue stars denote values of $t$ corresponding to the experiments in panels b, $t=2.75\text{\AA }$, c, $t=4.5\text{\AA }$ (also shown in Fig. A4c), and d, $t=9\text{\AA }$. Each method drew 100 samples per iteration. For each iteration, we show the mean (solid line) and 95% confidence interval (shaded envelope) of the 100 samples evaluated on $f(x)$, averaged across results from 20 random seeds that dictated initialization of the search distribution. P-values are from two-sided paired t-tests that AUC of the per-iteration mean is different between methods, using the 20 mean curves.